### v1

# Thermodynamic Entropic Uncertainty Relation

Abstract
Thermodynamic uncertainty relations reveal a fundamental trade-off between the precision of trajectory observables and entropy production, where the uncertainty in observables is quantified by their variance. In the context of information theory, uncertainty is often evaluated in terms of Shannon entropy, but it remains unclear whether there is a quantitative relation between Shannon entropy of the observables and entropy production in stochastic thermodynamics. In this Letter, we show that uncertainty relations can be formulated with observable Shannon entropy and entropy production. We introduce symmetry entropy, an entropy measure that quantifies the symmetry of the observable distribution, and demonstrate that a greater asymmetry in the observable distribution demands higher entropy production. This Letter elucidates the role of Shannon entropy of observables within stochastic thermodynamics, thereby establishing a foundation for deriving new uncertainty relations.
Introduction.—Quantum mechanics operates in ways that are fundamentally different from classical physics. The Heisenberg uncertainty relation, proposed by Heisenberg in 1927 [1], captures the unique nature of quantum mechanics through a single inequality, rendering the inability to determine position and momentum precisely. Robertson [2] generalized the Heisenberg uncertainty relation so that the relation can incorporate observables other than position or momentum. Given a quantum state |$\ket{\psi}$ and observables $A$ and $B$, the Robertson uncertainty relation states
$\mathrm{Var}[A]\mathrm{Var}[B]\geq\frac{1}{4}|\braket{\psi}{[A,B]}{\psi}|^{2},$ (1)
where [,][$\bullet$,$\bullet$][ ∙ , ∙ ] is the commutator and $\mathrm{Var}[A]$ is the variance of $A$ with respect to |$\ket{\psi}$. The Robertson uncertainty relation is recognized for its fundamental role in deriving various other relations, including the quantum speed limit [3, 4, 5]. Although the Robertson uncertainty relation given by Eq. (1) quantifies the uncertainty of the observables via their variance, uncertainty is often evaluated using the Shannon entropy in the context of information theory. It is therefore natural to expect uncertainty relations that incorporate the Shannon entropy. Indeed, Ref. [6] showed that the uncertainty relation involving the entropy of position and momentum holds in quantum mechanics. Several generalizations and extensions have been proposed for the entropic uncertainty relation [7] and the most well known instance is Maassen-Uffink relation [8, 9]:
$H[A]+H[B]\geq\ln\frac{1}{c}.$ (2)
Here, $H[A]$ and $H[B]$ denotes the Shannon entropy of measurement outputs of $A$ and $B$, respectively, and $c$ denotes the maximum overlap between two eigenvectors of $A$ and $B$. The entropic uncertainty relation is crucial not only for understanding the nature of quantum mechanics but also plays a vital role in quantum cryptography, especially in quantum key distribution protocols.
In recent years, it has become clear that uncertainty relations are prevalent in stochastic thermodynamics [10, 11]. In particular, the thermodynamic uncertainty relation [12, 13] indicates a trade-off between entropy production and the relative variance of the current observable. For a stochastic thermodynamic system in the steady state and a trajectory observable $F$, the following relation holds:
$\frac{\mathrm{Var}[F]}{\mathbb{E}[F]^{2}}\geq\frac{2}{\Sigma},$ (3)
where $\mathbb{E}[F]$ and $\mathrm{Var}[F]$ denote the expectation and the variance of $F$, respectively, and $\Sigma$ is the entropy production within the time interval of interest. Equation (3) suggests that achieving greater precision requires increased entropy production and signifies a no-free lunch in thermodynamic systems. The thermodynamic uncertainty relation in Eq. (3) resembles the Robertson uncertainty relation in that the uncertainty is evaluated with the variance. In fact, certain types of thermodynamic uncertainty relations are known to actually be derived from the Robertson uncertainty relation [14, 15]. Given this context, a simple question arises: Does an entropic uncertainty relation hold in stochastic thermodynamics? In this paper, we confirm that this conjecture is correct and demonstrate that there is an uncertainty relation between entropy production and the symmetry entropy [cf. Eq. (12)], which quantifies the extent of symmetry of observable distributions. Specifically, we show that the sum of the entropy production and the symmetry entropy should be no less than $\ln 2$ [cf. Eqs. (14) and (16)]. In other words, the asymmetry of the observable probability distribution, as quantified by entropy, requires that the entropy production be at least equal to this measure of asymmetry. We validate the uncertainty relation with a biased random walk, which is often used to model biomolecular processes. The trade-off relationship in thermodynamic cost has been extensively studied, particularly in terms of thermodynamic uncertainty relations [12, 13, 16, 17, 18, 19, 20, 21, 22, 23] using the variance of observable quantities and the speed limits using the distance between states [24, 25, 26, 22, 27, 28, 29]. This study demonstrates a trade-off between the entropy of observables and the entropy production, which is expected to lead to the derivation of other trade-off relations.
Figure 1: Examples of the symmetry entropy $\Lambda[P(\Phi)]$. Horizontal axes denote values of $\Phi$ (left column) and $|\Phi|$ (right column). The values of $\Phi$ always form pairs; $-1$ and $1$, and $-2$ and $2$ are two pairs in the examples. Vertical axes denote probability distribution $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column). (a) Symmetric distribution where probabilities of all of the pairs are identical. $P(|\Phi|)$ is different from the original distribution $P(\Phi)$ and thus $\Lambda[P(\Phi)]$ is $\ln 2$. (b) Asymmetric distribution where probabilities of all of the pairs are fully biased. $P(|\Phi|)$ and $P(\Phi)$ are identical and thus $\Lambda[P(\Phi)]$ is 00. (c) Asymmetric distribution where probabilities of all of the pairs are fully biased. The distributions $P(|\Phi|)$ and $P(\Phi)$ are different, but they effectively become the same if the labels for the pair $-1$ and $1$ are swapped. Therefore, they are essentially the same distribution. This results in $\Lambda[P(\Phi)]=0$.
Methods.—Let $X$ be a random variable and $P(X)$ be its probability distribution. The Shannon entropy $H[P(X)]$ is defined by
$H[P(X)]\equiv-\sum_{x}P(X=x)\ln P(X=x).$ (4)
Let $Y$ be another random variable. The Kullback-Leibler divergence between $P(X)$ and $P(Y)$ is defined by
$D[P(X)\|P(Y)]$ $\equiv\sum_{x}P(X=x)\ln\frac{P(X=x)}{P(Y=x)}$
$=-H[P(X)]+C[P(X),P(Y)].$ (5)
where $C[P(X),P(Y)]$ is the cross entropy:
$C[P(X),P(Y)]\equiv-\sum_{x}P(X=x)\ln P(Y=x).$ (6)
It is known that the Kullback-Leibler divergence is non-negative. Moreover, the Kullback-Leibler divergence satisfies monotonicity. Consider a transformation that maps the original random variables $X$ and $Y$ to new random variables $\tilde{X}$ and $\tilde{Y}$, respectively. Then the following monotonicity relation holds:
$D[P(X)\|P(Y)]\geq D[P(\tilde{X})\|P(\tilde{Y})],$ (7)
where $P(\tilde{X})$ and $P(\tilde{Y})$ are probability distributions of the transformed variables $\tilde{X}$ and $\tilde{Y}$. The Kullback-Leibler divergence is a measure that quantifies the distance between two probability distributions. The monotonicity shown by Eq. (7) implies that no matter what transformation is applied to random variables, the distance between the two probability distributions cannot be increased.
Having introduced basic concepts of the divergence, we move to consideration of stochastic thermodynamic systems. Stochastic thermodynamics considers processes whose state change is described by a stochastic process. Let $\Gamma$ be a stochastic trajectory of the process and $\Gamma^{\dagger}$ be its time reversal. Moreover, we can define the probability of measuring $\Gamma$, which is denoted by $\mathcal{P}(\Gamma)$. Assuming the local detailed balance, it is known that the entropy production under the steady-state condition is defined by the Kullback-Leibler divergence:
$\Sigma=D[\mathcal{P}(\Gamma)\|\mathcal{P}(\Gamma^{\dagger})].$ (8)
Note that the expression of Eq. (8) also holds for continuous processes such as Langevin dynamics. Consider an observable $\Phi(\Gamma)$, which is a function of the trajectory $\Gamma$. Here, we assume that $\Phi(\Gamma)$ is anti-symmetric under the time reversal:
$\Phi(\Gamma)=-\Phi(\Gamma^{\dagger}).$ (9)
For example, $\Phi(\Gamma)$ represents a thermodynamic current. Important thermodynamic quantities, such as stochastic dissipated heat or displacement, are expressed by $\Phi(\Gamma)$. Here, we initially assume that $\Phi(\Gamma)$ takes values in a countable set, that is, the probability distribution $P(\Phi)$ is discrete. However, all the results below hold for the continuous case as well.
Figure 2: Illustration of the example. (a) Biased random walk with right and left transition probabilities are $w_{+}$ and $w_{-}$, respectively. (b) Illustration of F1-ATPase. This biased random walk can model molecular motors such as F1-ATPase, where the motor uses ATP hydrolysis to rotate discretely in steps of $32\pi/$ 3 in the counter-clockwise direction.
Results.—We derive the thermodynamic entropic uncertainty relation, which is the main result of this Letter. Using the monotonicity of Kullback-Leibler divergence [Eq. (7)] and the time reversal property of $\Phi$ [Eq. (9)], we have
$\Sigma$ $\geq D[P(\Phi)\|P(-\Phi)]$
$=-H[P(\Phi)]+C[P(\Phi),P(-\Phi)].$ (10)
Here, the cross entropy is evaluated as $C[P(\Phi),P(-\Phi)]=-\sum_{\phi}P(\Phi=\phi)\ln P(\Phi=-\phi)$. Since the cross entropy term is no less than the entropy, it is non-negative, $C[P(\Phi),P(-\Phi)]\geq H[P(\Phi)]\geq 0$. An important observable is that $C[P(\Phi),P(-\Phi)]$ is even bounded from below by a positive term. To simplify, let us consider that the observable $\Phi(\Gamma)$ does not include 00, whose condition is met for several problem settings. For instance, we may consider binary classification using trajectories of stochasitc processes. In this case, the observable $\Phi(\Gamma)$ does not include 00. $\Phi(\Gamma)=0$ should be handled separately because $\Phi(\Gamma)=\Phi(\Gamma^{\dagger})=0$ from Eq. (9), showing that the observable is invariant under the time reversal. Later, we will consider the case where $\Phi(\Gamma)=0$ is included. Specifically, it can be proved that the following relation holds:
$C[P(\Phi),P(-\Phi)]\geq\ln 2+H[P(|\Phi|)],$ (11)
whose derivation is shown in Appendix A. In Eq. (11), $P(|\Phi|)$ is the probability distribution of $|\Phi|$, which is the absolute random variable of $\Phi$. Let $\Lambda[P(\Phi)]$ be
$\Lambda[P(\Phi)]\equiv H[P(\Phi)]-H[P(|\Phi|)],$ (12)
which is the entropy difference between the original distribution $P(\Phi)$ and its absolute valued distribution $P(|\Phi|)$. Here, we call $\Lambda[P(\Phi)]$ as symmetry entropy. The symmetry entropy $\Lambda[P(\Phi)]$ quantifies the extent of symmetry of $P(\Phi)$. Figure 1 depicts examples of values of the symmetry entropy, where the horizontal axes are values of $\Phi$ (left column) and $|\Phi|$ (right column) and the vertical axes are $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column). From the condition of time reversal [Eq. (9)], if $\Phi=a$ exists ($a>0$), then $\Phi=-a$ also exists, which is regarded as a pair of the observable. Suppose that the distribution $P(\Phi)$ is symmetric for all of the pairs, as depicted in Fig. 1(a). For this case, $P(|\Phi|)$ is very different from the original distribution $P(\Phi)$, which results in $\Lambda[P(\Phi)]$ being $\ln 2$. In contrast, for an asymmetric distribution illustrated in Fig. 1(b), $P(\Phi)$ and $P(|\Phi|)$ are the same, and thus $\Lambda[P(\Phi)]$ reduces to 00. To be more specific, the symmetry entropy quantifies how biased the probability distributions of pairs of observables are. When there is a strong bias in the probabilities of the pairs, the value of $\Lambda[P(\Phi)]$ becomes small. For example, $\Lambda[P(\Phi)]$ becomes 00 not only in cases where $P(\Phi)$ and $P(|\Phi|)$ are the same, as shown in Fig. 1(b), but also in cases like Fig. 1(c). It can be shown that
$0\leq\Lambda[P(\Phi)]\leq\ln 2,$ (13)
whose proof is provided in Appendix B. In Eq. (13), $\Lambda[P(\Phi)]$ being 00 and $\ln 2$ corresponds to asymmetric and symmetric distributions, respectively. Substituting Eq. (11) into Eq. (10), we obtain the trade-off between the entropy production and the asymmetry of $P(\Phi)$ quantified by entropy:
$\Sigma\geq\ln 2-\Lambda[P(\Phi)]\geq 0.$ (14)
Equation (14) is the main result of this study and referred to as thermodynamic entropic uncertainty relation. The right-hand side of Eq. (14) quantifies the asymmetry of probability distribution $P(\Phi)$. Therefore, Eq. (14) shows that, for arbitrary observable $\Phi$ satisfying the time reversal condition [Eq. (9)], the system requires the entropy production no less than $\ln 2-\Lambda[P(\Phi)]$. The trade-off between entropy production and observable asymmetry parallels traditional thermodynamic uncertainty relations, which illustrate a trade-off between the variance of observables and entropy production. The thermodynamic uncertainty relations consider the relative variance $\mathrm{Var}[\Phi]/\mathbb{E}[\Phi]^{2}$. In a sense, using the relative variance can also be seen as quantifying the asymmetry of $P(\Phi)$; when the variance is smaller and the expectation is greater, the probability distribution $P(\Phi)$ is more asymmetric with respect to $\Phi=0$. Since $\ln 2-\Lambda[P(\Phi)]\geq 0$ in Eq. (14) due to Eq. (13), Eq. (14) can be regarded as a refinement of the second law using the entropy of the observable. There are some advantages in employing the entropy instead of the variance. When dealing with the variance, the observable must yield a real number. However, in cases where the observable consists of classifications such as “success” and “failure”, the variance is not suitable. Even in such cases, the Shannon entropy is well defined, indicating that Eq. (14) can be applied.
Thus far, we assumed that $\Phi(\Gamma)$ takes values in a countable set, that is, $P(\Phi)$ is a discrete distribution. Consider the case where $\Phi(\Gamma)$ produces continuous values, where the summation should be replaced by the integration. When considering a continuous distribution, the notable difference is that the differential entropy may take negative values. However, the existence of negative values is not problematic, as such negative values are offset in $H[P(\Phi)]-H[P(|\Phi|)]$. Therefore, the main result of Eq. (14) is well defined for the continuous case as well.
The derivation of Eq. (14) assumed that the observable $\Phi(\Gamma)$ does not include $\Phi(\Gamma)=0$. It is straightforward to extend the result to the case where the observable includes $\Phi(\Gamma)=0$. Specifically, when $\Phi(\Gamma)=0$ is included, the range of $\Lambda[P(\Phi)]$ is modified as follows:
$0\leq\Lambda[P(\Phi)]\leq[1-P(\Phi=0)]\ln 2$ (15)
The bound becomes
$\Sigma\geq[1-P(\Phi=0)]\ln 2-\Lambda[P(\Phi)]\geq 0,$ (16)
which includes Eq. (14) as the specific case $P(\Phi=0)=0$. As long as the entropy production is given by Eq. (8), Eq. (16) holds for arbitrary observable $\Phi(\Gamma)$ satisfying Eq. (9). In Eq. (16), $\Phi=0$ plays a special role. When considering the process of doing nothing, the observable $\Phi$ is always 00 implying $P(\Phi=0)=1$. For such empty dynamics, the entropy production is 00, showing that both sides of Eq. (16) equal 00 and thus the inequality becomes equality. This consideration shows that $\Phi=0$ should be handled separately. However, note that this exceptional handling of $\Phi=0$ is not limited to Eq. (16). In the conventional thermodynamic uncertainty relation [Eq. (3)], when the expectation of the current vanishes, the inequality is ill-defined. When $P(\Phi)$ is a probability density that is smooth around 00, the measure becomes $P(\Phi=0)=0$. Only when $P(\Phi)$ includes the delta-peaked contribution at $\Phi=0$, we use Eq. (16)
Let us comment on the relation between Eq. (16) and the Landauer principle [30]. When considering a process that resets to one state from a state that exists with equal probability in two states, the following relation is obtained.
$\Delta S_{m}\geq\ln 2,$ (17)
where $\Delta S_{m}$ is the entropy increase in the surrounding medium. Note that the Landauer principle can be considered as a specific case of the second law of thermodynamics. Given the assumption that the system is in steady state, the change in system entropy $\Delta S$ is zero; consequently, $\Sigma=\Delta S_{m}$, where $\Delta S_{m}$ represents the increase in entropy in the environment. Then, Eq. (14) can be expressed by $\Delta S_{m}\geq\ln 2-\Lambda[P(\Phi)]$. When the probability distribution of $P(\Phi)$ ends up with a totally asymmetric distribution, i.e., $\Lambda[P(\Phi)]=0$, Eq. (14) is formally identical to the Landauer principle. However, note that the right side of Eq. (17) arises from a reduction in the Shannon entropy within the system’s state, implying that the Landauer principle is relevant when the system state changes over the time evolution. This contrast with Eq. (14), where the system is assumed to be steady state, and $\ln 2$ on its right-hand side arises from the asymmetry of the observable, not the system state.
Example.—To illustrate Eq. (14), consider a one-dimensional biased random walk as shown in Fig. 2(a). We consider a discrete time model, where the transition probabilities for moving right and left are denoted as $w_{+}$ and $w_{-}$, respectively, with the constraint $w_{+}+w_{-}=1$. The biased random walk plays an important role in biological modeling. For example, a biased random walk can model molecular motors, such as F1-ATPase [31] (Fig. 2(b)). F1-ATPase is an enzyme that can rotate in discrete steps of $32\pi/$ 3 radians in the clockwise or counterclockwise direction [32]. When rotating counterclockwise, F1-ATPase hydrolyzes ATP, while when rotating clockwise, F1-ATPase synthesizes ATP. The rotation in both directions is stochastic, which means that it occurs randomly. The counter-clockwise rotation can be considered as moving in the positive direction in the biased random walk, where the motor has a higher probability of rotating in the counter-clockwise direction.
In the biased random walk, we consider the observable $\Phi(\Gamma)$, which takes either “$+$” or “--”. After the $L$ steps, the observable $\Phi$ is determined by comparing the final position to the initial position. If the final position is greater than the initial position, $\Phi$ is assigned the value “$+$”. Conversely, if the final position is less than or equal to the initial position, $\Phi$ is assigned the value “--”. When the number of steps is odd, there is no possibility of $\Phi(\Gamma)=0$. When considering the aforementioned molecular motor, this observable corresponds to checking whether the motor rotates in the counterclockwise direction or not. For details, see Appendix C.
Figure 3: Results of the biased random walk. The sum of the entropy production and the symmetry entropy, $\Sigma+\Lambda[P(\Phi)]$, is plotted as a function of $w_{+}$ with the solid line. The dotted line denotes $\ln 2$, which is the lower bound of $\Sigma+\Lambda[P(\Phi)]$ in Eq. (14). Each of these terms, $\Sigma$ and $\Lambda[P(\Phi)]$, is plotted by the dashed line and the dot-dashed line, respectively. Here, we employ $L=5$.
We calculate quantities, such as $\Sigma$ and $\Lambda[P(\Phi)]$, as a function of $w_{+}$ in Fig. 3. Note that, for the binary case where $\Phi(\Gamma)$ takes two values, $\Lambda[P(\Phi)]=H[P(\Phi)]$. We consider the steady-state condition and calculate the probability distribution $P(\Phi)$ and $\Sigma$. In Fig. 3, $\Sigma+\Lambda[P(\Phi)]$, the sum of the entropy production $\Sigma$ and the symmetry entropy $\Lambda[P(\Phi)]$, is represented by the solid line, where the dotted line indicates $\ln 2$. As can be seen in the figure, $\Sigma+\Lambda[P(\Phi)]$ is always above $\ln 2$, confirming the validity of Eq. (14). To examine how each of these components, $\Sigma$ and $\Lambda[P(\Phi)]$, behave as a function of $w_{+}$, $\Sigma$ and $\Lambda[P(\Phi)]$ are represented by the dashed line and the dot-dashed line, respectively. Regarding entropy production $\Sigma$, it falls below $\ln 2$ in the region where $w_{+}$ ranges from $0.5$ to $0.65$. On the other hand, the symmetry entropy of the observable $\Lambda[P(\Phi)]$ is decreasing as a function of $w_{+}$, and except for $w_{+}=0.5$, it is below $\ln 2$. From these results, we can conclude that the inequality $\Sigma+\Lambda[P(\Phi)]\geq\ln 2$ holds only when considering the sum, $\Sigma+\Lambda[P(\Phi)]$.
Conclusion.—In this study, we established a thermodynamic entropic uncertainty relation that links entropy production and the Shannon entropy of the observables. Our findings extend conventional thermodynamic uncertainty relations by incorporating measures based on entropy, highlighting the role of Shannon entropy of observables in stochastic thermodynamics. The derived inequality formalizes a fundamental trade-off between entropy production and the asymmetry of the observable distribution. We validated this relation using a biased random walk model. This framework provides a deeper understanding of nonequilibrium thermodynamics and expands the application of entropy-based uncertainty relations in stochastic systems. One direction of expansion is towards quantum systems. In recent years, the thermodynamic uncertainty relations in quantum systems [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 14, 43] have garnered significant attention. In particular, the thermodynamic uncertainty relations within the framework of continuous measurement are closely related to those in classical stochastic processes. This direction presents future challenges.
Appendix A Derivation of the main result [Eq. (16)]
A.1 Discrete case
In this section, we derive Eq. (16). Let us introduce the absolute random variable of $\Phi$, which is expressed by $|\Phi|$. The probability distribution of $|\Phi|$ is given by
$P(|\Phi|=\phi)=\begin{cases}P(\Phi=\phi)+P(\Phi=-\phi)&\phi\neq 0\\ P(\Phi=0)&\phi=0\end{cases}.$ (18)
Therefore, the Shannon entropy of $|\Phi|$ is given by
$H[P(|\Phi|)]$
$=-\sum_{\phi\geq 0}P(|\Phi|=\phi)\ln P(|\Phi|=\phi)$
$=-P(0)\ln P(0)-\sum_{\phi>0}[P(\phi)+P(-\phi)]\ln[P(\phi)+P(-\phi)]$
$=-P(0)\ln P(0)-\sum_{\phi>0}P(\phi)\ln[P(\phi)+P(-\phi)]$
$-\sum_{\phi<0}P(\phi)\ln[P(\phi)+P(-\phi)]$
$=-\sum_{\phi}P(\phi)\ln[P(\phi)+P(-\phi)]+P(0)\ln 2,$ (19)
where we abbreviated $P(\phi)=P(\Phi=\phi)$. Then we compute $H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$ as follows:
$H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$
$=\sum_{\phi}P(\phi)\ln\frac{P(-\phi)/P(\phi)}{1+P(-\phi)/P(\phi)}+P(0)\ln 2.$ (20)
Let us consider the function $f(x)=\ln\frac{x}{1+x}$. Since $f(x)$ is concave for $x>0$, by using the Jensen inequality, the following relation holds:
$\sum_{\phi}P(\phi)\ln\frac{P(-\phi)/P(\phi)}{1+P(-\phi)/P(\phi)}$
$\leq\ln\frac{\sum_{\phi}P(\phi)\left(P(-\phi)/P(\phi)\right)}{1+\sum_{\phi}P(\phi)\left(P(-\phi)/P(\phi)\right)}$
$=-\ln 2.$ (21)
By substituting Eq. (21) into Eq. (20), we obtain
$H[P(|\Phi|)]+[1-P(\Phi=0)]\ln 2$
$\leq C[P(\Phi),P(-\Phi)].$ (22)
Equation (22) proves Eq. (11) in the main text.
A.2 Continuous case
Next, we consider the continuous case. Basically, the derivation is the same as in the discrete case except that the summation should be replaced by the integration.
Assume $P(\Phi)$ represents a probability density that is differentiable for all $\Phi$. For the continuous case, the probability density $P(|\Phi|)$ is defined by
$P(|\Phi|=\phi)=P(\Phi=\phi)+P(\Phi=-\phi).$ (23)
We do not consider $P(|\Phi|=0)$, because such measure is 00 for the smooth probability density. When $P(\Phi)$ includes the contribution of the delta function in $\Phi=0$, this is not the case. The continuous entropy is given by
$H[P(|\Phi|)]$
[…]
Again, we calculate $H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$ as follows:
$H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$
[…]
Using Eq. (25), we prove
$H[P(|\Phi|)]+\ln 2$ $\leq C[P(\Phi),P(-\Phi)].$ (26)
Appendix B Proof of Eq. (13)
In this section, we prove Eq. (13). Here, we show the relation for $P(\Phi=0)\geq 0$. The relation which we want to show is given by
$0\leq\Lambda[P(\Phi)]\leq[1-P(\Phi=0)]\ln 2$ (27)
The first inequality part corresponds to $H[P(|\Phi|)]\leq H[P(\Phi)]$. For random variables $X$ and $Y$, the following relation holds:
$H[P(X),P(Y)]$ $=H[P(X)|P(Y)]+H[P(Y)]$
$=H[P(Y)|P(X)]+H[P(X)].$ (28)
Substituting $X=\Phi$ and $Y=|\Phi|$, we obtain
$H[P(|\Phi|)|P(\Phi)]+H[P(\Phi)]$ $=H[P(\Phi)]$
$=H[P(\Phi)|P(|\Phi|)]+H[P(|\Phi|)]$
$\geq H[P(|\Phi|)],$ (29)
where we used $H[P(|\Phi|)|P(\Phi)]=0$. Equation (29) proves the first inequality part of Eq. (27). Note that, for the continuous case, this proof does not work, as the continuous conditional entropy $H[P(\Phi)|P(|\Phi|)]$ may take negative values. However, we can show that the first inequality part also holds for the continuous case by a direct calculation.
Next, we prove the second part of the inequality of Eq. (27), which can be done following the same approach as in Appendix A. Using Eq. (19), we have
$H[P(|\Phi|)]-H[P(\Phi)]$
$=-\sum_{\phi}P(\phi)\ln\left(1+\frac{P(-\phi)}{P(\phi)}\right)+P(0)\ln 2$
$\geq-\ln 2+P(0)\ln 2,$ (30)
where we again used the Jensen inequality. Equation (30) proves the second inequality part of Eq. (27). For the continuous case, by following the same argument by replacing the summation with the integration, the following relation holds:
$H[P(|\Phi|)]-H[P(\Phi)]$ $\geq-\ln 2.$ (31)
Appendix C Details of example
We show details of the biased random walk example. The probability distribution at position $x$ after $L$ steps is given by
$p(x)=\begin{cases}{\displaystyle\binom{L}{\frac{L+x}{2}}p^{\frac{L+x}{2}}(1-p)^{\frac{L-x}{2}}}&L+x\in\mathrm{even}\\ 0&\mathrm{otherwise}\end{cases}.$ (32)
Using $p(x)$ in Eq. (32), for odd $L$, the probability of $\Phi$ is calculated by $P(\Phi=\mathrm{decrease})=\sum_{x=-L}^{-1}p(x)$ and $P(\Phi=\mathrm{increase})=\sum_{x=1}^{L}p(x)$, . which can be used to calculate the Shannon entropy of $P(\Phi)$. Moreover, under the steady-state condition, the entropy production $\Sigma$ up to $L$ steps is given by
$\Sigma=L(w_{+}-w_{-})\ln\frac{w_{+}}{w_{-}}.$ (33)

### v3

@@
 Abstract
-Thermodynamic uncertainty relations reveal a fundamental trade-off between the precision of trajectory observables and entropy production, where the uncertainty in observables is quantified by their variance.
-In the context of information theory, uncertainty is often evaluated in terms of Shannon entropy, but it remains unclear whether there is a quantitative relation between Shannon entropy of the observables and entropy production in stochastic thermodynamics.
-In this Letter, we show that uncertainty relations can be formulated with observable Shannon entropy and entropy production.
-We introduce symmetry entropy, an entropy measure that quantifies the symmetry of the observable distribution, and demonstrate that a greater asymmetry in the observable distribution demands higher entropy production.
-This Letter elucidates the role of Shannon entropy of observables within stochastic thermodynamics, thereby establishing a foundation for deriving new uncertainty relations.
-Introduction.—Quantum mechanics operates in ways that are fundamentally different from classical physics.
-The Heisenberg uncertainty relation, proposed by Heisenberg in 1927 [1], captures the unique nature of quantum mechanics through a single inequality, rendering the inability to determine position and momentum precisely.
-Robertson [2] generalized the Heisenberg uncertainty relation so that the relation can incorporate observables other than position or momentum.
-Given a quantum state |$\ket{\psi}$ and observables $A$ and $B$, the Robertson uncertainty relation states
-$\mathrm{Var}[A]\mathrm{Var}[B]\geq\frac{1}{4}|\braket{\psi}{[A,B]}{\psi}|^{2},$ (1)
-where [,][$\bullet$,$\bullet$][ ∙ , ∙ ] is the commutator and $\mathrm{Var}[A]$ is the variance of $A$ with respect to |$\ket{\psi}$.
-The Robertson uncertainty relation is recognized for its fundamental role in deriving various other relations, including the quantum speed limit [3, 4, 5].
-Although the Robertson uncertainty relation given by Eq.
-(1) quantifies the uncertainty of the observables via their variance, uncertainty is often evaluated using the Shannon entropy in the context of information theory.
-It is therefore natural to expect uncertainty relations that incorporate the Shannon entropy.
-Indeed, Ref.
-[6] showed that the uncertainty relation involving the entropy of position and momentum holds in quantum mechanics.
-Several generalizations and extensions have been proposed for the entropic uncertainty relation [7] and the most well known instance is Maassen-Uffink relation [8, 9]:
-$H[A]+H[B]\geq\ln\frac{1}{c}.$ (2)
-Here, $H[A]$ and $H[B]$ denotes the Shannon entropy of measurement outputs of $A$ and $B$, respectively, and $c$ denotes the maximum overlap between two eigenvectors of $A$ and $B$.
-The entropic uncertainty relation is crucial not only for understanding the nature of quantum mechanics but also plays a vital role in quantum cryptography, especially in quantum key distribution protocols.
-In recent years, it has become clear that uncertainty relations are prevalent in stochastic thermodynamics [10, 11].
-In particular, the thermodynamic uncertainty relation [12, 13] indicates a trade-off between entropy production and the relative variance of the current observable.
-For a stochastic thermodynamic system in the steady state and a trajectory observable $F$, the following relation holds:
+Thermodynamic uncertainty relations reveal a fundamental trade-off between the precision of a trajectory observable and entropy production, where the uncertainty of the observable is quantified by its variance.
+In information theory, Shannon entropy is a common measure of uncertainty.
+However, a clear quantitative relationship between the Shannon entropy of an observable and the entropy production in stochastic thermodynamics remains to be established.
+In this Letter, we show that an uncertainty relation can be formulated in terms of the Shannon entropy of an observable and the entropy production.
+We introduce symmetry entropy, an entropy measure that quantifies the symmetry of the observable distribution, and demonstrate that a greater asymmetry in the observable distribution requires higher entropy production.
+Specifically, we establish that the sum of the entropy production and the symmetry entropy cannot be less than $\ln 2$.
+As a corollary, we also prove that the sum of the entropy production and the Shannon entropy of the observable is no less than $\ln 2$.
+As an application, we demonstrate our relation in the diffusion decision model, revealing a fundamental trade-off between decision accuracy and entropy production in stochastic decision-making processes.
+Introduction.— The thermodynamic uncertainty relation [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] expresses a universal trade-off between the entropy production and the relative variance of time-integrated observables for systems operating far from equilibrium.
+For a steady-state stochastic thermodynamic system and a trajectory observable $F$, the thermodynamic uncertainty relation states [1, 2]
 $\frac{\mathrm{Var}[F]}{\mathbb{E}[F]^{2}}\geq\frac{2}{\Sigma},$ (1)
-where $\mathbb{E}[F]$ and $\mathrm{Var}[F]$ denote the expectation and the variance of $F$, respectively, and $\Sigma$ is the entropy production within the time interval of interest.
-Equation (3) suggests that achieving greater precision requires increased entropy production and signifies a no-free lunch in thermodynamic systems.
-The thermodynamic uncertainty relation in Eq.
-(3) resembles the Robertson uncertainty relation in that the uncertainty is evaluated with the variance.
-In fact, certain types of thermodynamic uncertainty relations are known to actually be derived from the Robertson uncertainty relation [14, 15].
-Given this context, a simple question arises: Does an entropic uncertainty relation hold in stochastic thermodynamics?
-In this paper, we confirm that this conjecture is correct and demonstrate that there is an uncertainty relation between entropy production and the symmetry entropy [cf.
+where $\mathbb{E}[F]$ and $\mathrm{Var}[F]$ are the expectation value and variance of $F$, respectively, and $\Sigma$ denotes the entropy production during the observation interval.
+The thermodynamic uncertainty relation implies that achieving higher precision inevitably requires greater entropy production–a manifestation of the “no free lunch” principle in thermodynamics.
+The thermodynamic uncertainty relation given by Eq.
+(1) quantifies the uncertainty of the observables via their variance; however, uncertainty is often evaluated using the Shannon entropy in the context of information theory.
+Consider a thermodynamic decision-making process that yields a categorical output, such as a “yes” or “no” label.
+Since the output is categorical rather than numerical, we cannot use variance to quantify its uncertainty.
+However, even in such cases, we can still calculate the Shannon entropy.
+Therefore, to analyze the tradeoff between the accuracy of the categorical outputs and the entropy production, a thermodynamic uncertainty relation based on Shannon entropy is required.
+The issue of how to quantify uncertainty has been actively studied in quantum mechanics.
+In the Heisenberg-Robertson uncertainty relation [11, 12], the uncertainty in the outcomes of two observables is quantified through their variances.
+Later, Hirschman showed that the variance can be replaced by the Shannon entropy [13, 14].
+Maassen and Uffink [15] derived a precise statement of this entropic uncertainty relation [16]: the sum of the entropies of the two observables cannot fall below a bound determined by the extent of their eigenstate overlap.
+The entropic uncertainty relation plays a fundamental role in quantum cryptography, especially in quantum key distribution protocols [17].
+In this Letter, in the context of stochastic thermodynamics, we show that there is an uncertainty relation between entropy production and the symmetry entropy [cf.
 Eq.
-(12)], which quantifies the extent of symmetry of observable distributions.
-Specifically, we show that the sum of the entropy production and the symmetry entropy should be no less than $\ln 2$ [cf.
+(8)], which quantifies the extent of symmetry of the observable’s distribution.
+Specifically, we show that the sum of the entropy production and the symmetry entropy of the observable should be no less than $\ln 2$ [cf.
 Eqs.
@@
 In other words, the asymmetry of the observable probability distribution, as quantified by entropy, requires that the entropy production be at least equal to this measure of asymmetry.
-We validate the uncertainty relation with a biased random walk, which is often used to model biomolecular processes.
-The trade-off relationship in thermodynamic cost has been extensively studied, particularly in terms of thermodynamic uncertainty relations [12, 13, 16, 17, 18, 19, 20, 21, 22, 23] using the variance of observable quantities and the speed limits using the distance between states [24, 25, 26, 22, 27, 28, 29].
-This study demonstrates a trade-off between the entropy of observables and the entropy production, which is expected to lead to the derivation of other trade-off relations.
+As a corollary of the result, we also show that the sum of the entropy production and the Shannon entropy of the observable should be no less than $\ln 2$ [Eq.
+(12)].
+We apply the derived relation to the diffusion decision model, illustrating a trade-off between decision accuracy and entropy production in stochastic decision-making processes.
+The trade-off relation in thermodynamic cost has been actively studied, particularly in terms of thermodynamic uncertainty relations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] using the variance of observable quantities and speed limits using the distance between states [18, 19, 20, 9, 21, 22, 23].
+This study demonstrates a trade-off between the entropy of the observable and the entropy production, which is expected to lead to the derivation of other trade-off relations.
+Figure 1: Conceptual representation of the thermodynamic entropic uncertainty relation.
+(a) Stochastic thermodynamic process.
+The thermodynamic entropic uncertainty relation considers a stochastic process, where each state transition is random.
+(b) Trajectory of the stochastic process shown in (a).
+$\Gamma$ denotes time evolution of a realization of the process.
+$\Gamma^{\dagger}$ is the time-reversal of $\Gamma$.
+(c) Probability distribution of observable $\Phi(\Gamma)$.
+$\Phi(\Gamma)$ is arbitrary, provided it satisfies the time-reversal antisymmetric property [Eq.
+(7)].
 Figure 2: Examples of the symmetry entropy $\Lambda[P(\Phi)]$.
@@
 Vertical axes denote probability distribution $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column).
-(a) Symmetric distribution where probabilities of all of the pairs are identical.
+(a) Distribution symmetric with respect to $\Phi=0$ where probabilities of all of the pairs are identical.
 $P(|\Phi|)$ is different from the original distribution $P(\Phi)$ and thus $\Lambda[P(\Phi)]$ is $\ln 2$.
@@
 $C[P(X),P(Y)]\equiv-\sum_{x}P(X=x)\ln P(Y=x).$ (4)
-It is known that the Kullback-Leibler divergence is non-negative.
+It is well known that the Kullback-Leibler divergence is non-negative.
 Moreover, the Kullback-Leibler divergence satisfies monotonicity.
@@
 $D[P(X)\|P(Y)]\geq D[P(\tilde{X})\|P(\tilde{Y})],$ (5)
-where $P(\tilde{X})$ and $P(\tilde{Y})$ are probability distributions of the transformed variables $\tilde{X}$ and $\tilde{Y}$.
-The Kullback-Leibler divergence is a measure that quantifies the distance between two probability distributions.
-The monotonicity shown by Eq.
-(7) implies that no matter what transformation is applied to random variables, the distance between the two probability distributions cannot be increased.
-Having introduced basic concepts of the divergence, we move to consideration of stochastic thermodynamic systems.
-Stochastic thermodynamics considers processes whose state change is described by a stochastic process.
-Let $\Gamma$ be a stochastic trajectory of the process and $\Gamma^{\dagger}$ be its time reversal.
+where $P(\tilde{X})$ and $P(\tilde{Y})$ are probability distributions of the transformed variables $\tilde{X}$ and $\tilde{Y}$, respectively.
+Having introduced basic concepts of the divergence, we move to consider stochastic thermodynamic systems.
+Stochastic thermodynamics considers processes whose state change is described by a stochastic process (Fig.
+1(a)).
+Let $\Gamma$ be a stochastic trajectory of the process and $\Gamma^{\dagger}$ be its time reversal (Fig.
+1(b)).
 Moreover, we can define the probability of measuring $\Gamma$, which is denoted by $\mathcal{P}(\Gamma)$.
-Assuming the local detailed balance, it is known that the entropy production under the steady-state condition is defined by the Kullback-Leibler divergence:
+Assuming the local detailed balance, it is known that the entropy production under the steady-state condition is defined by the Kullback-Leibler divergence [24, 25]:
 $\Sigma=D[\mathcal{P}(\Gamma)\|\mathcal{P}(\Gamma^{\dagger})].$ (6)
@@
 Consider an observable $\Phi(\Gamma)$, which is a function of the trajectory $\Gamma$.
-Here, we assume that $\Phi(\Gamma)$ is anti-symmetric under the time reversal:
+Here, we assume that $\Phi(\Gamma)$ is antisymmetric under the time reversal:
 $\Phi(\Gamma)=-\Phi(\Gamma^{\dagger}).$ (7)
@@
 Important thermodynamic quantities, such as stochastic dissipated heat or displacement, are expressed by $\Phi(\Gamma)$.
-Here, we initially assume that $\Phi(\Gamma)$ takes values in a countable set, that is, the probability distribution $P(\Phi)$ is discrete.
-However, all the results below hold for the continuous case as well.
-Figure 2: Illustration of the example.
-(a) Biased random walk with right and left transition probabilities are $w_{+}$ and $w_{-}$, respectively.
-(b) Illustration of F1-ATPase.
-This biased random walk can model molecular motors such as F1-ATPase, where the motor uses ATP hydrolysis to rotate discretely in steps of $32\pi/$ 3 in the counter-clockwise direction.
+In the supplementary material [26], we provide a generalized formulation where the time-reversal antisymmetric condition [Eq.
+(7)] is relaxed.
+Initially, we assume that $\Phi(\Gamma)$ takes values in a countable set, that is, the probability distribution $P(\Phi)$ is discrete (Fig.
+1(c)).
+However, most of the results below hold for the continuous case as well.
+Figure 3: (a) Description of the diffusion decision model.
+The green trajectory illustrates a realization that exceeds the threshold $\theta$, resulting in a decision of $+1$ (“yes” class).
+Conversely, the purple trajectory represents a realization that falls below $-\theta$, leading to a decision of $-1$ (“no” class).
+(b) and (c) Entropies as a function of parameter $f$.
+In both figures, the blue dashed line shows the entropy production $\Sigma$, and the purple dot-dashed line represents $H[P(\Phi)]$, the entropy of the distribution $P(\Phi)$.
+Their sum, $\Sigma+H[P(\Phi)]$, is also depicted by the red solid line.
+The orange dotted line represents $\ln 2$ and $(1-P(0))\ln 2$ in (a) and (b), respectively, which serves as the lower bound for $\Sigma+H[P(\Phi)]$ according to Eq.
+(15).
+These plots are shown for $\theta=0$ in (b) and $\theta=0.5$ in (c), with other parameters being $A=1$ and $\tau=10$.
 Results.—We derive the thermodynamic entropic uncertainty relation, which is the main result of this Letter.
-Using the monotonicity of Kullback-Leibler divergence [Eq.
-(7)] and the time reversal property of $\Phi$ [Eq.
-(9)], we have
-$\Sigma$ $\geq D[P(\Phi)\|P(-\Phi)]$
-$=-H[P(\Phi)]+C[P(\Phi),P(-\Phi)].$ (10)
-Here, the cross entropy is evaluated as $C[P(\Phi),P(-\Phi)]=-\sum_{\phi}P(\Phi=\phi)\ln P(\Phi=-\phi)$.
-Since the cross entropy term is no less than the entropy, it is non-negative, $C[P(\Phi),P(-\Phi)]\geq H[P(\Phi)]\geq 0$.
-An important observable is that $C[P(\Phi),P(-\Phi)]$ is even bounded from below by a positive term.
-To simplify, let us consider that the observable $\Phi(\Gamma)$ does not include 00, whose condition is met for several problem settings.
-For instance, we may consider binary classification using trajectories of stochasitc processes.
+We first introduce an entropic measure which quantifies the observable of a trajectory $\Gamma$.
+Let $\Lambda[P(\Phi)]$ be
+$\Lambda[P(\Phi)]\equiv H[P(\Phi)]-H[P(|\Phi|)],$ (8)
+which is the entropy difference between the original distribution $P(\Phi)$ and its absolute valued distribution $P(|\Phi|)$.
+Here, we call $\Lambda[P(\Phi)]$ symmetry entropy.
+The symmetry entropy $\Lambda[P(\Phi)]$ quantifies the extent of symmetry of $P(\Phi)$.
+Intuitively speaking, the difference $H[P(\Phi)]-H[P(|\Phi|)]$ quantifies how much “sign” information remains in the distribution $P(\Phi)$ once its magnitude $P(|\Phi|)$ is known.
+If $P(\Phi)$ is perfectly symmetric with respect to $\Phi=0$, knowing $P(|\Phi|)$ tells us nothing about the sign, so $\Lambda[P(\Phi)]$ reaches its maximum of $\ln 2$ (one bit).
+In contrast, if $P(\Phi)$ is fully biased in one direction, the sign is already determined by $P(|\Phi|)$ and $\Lambda[P(\Phi)]$ vanishes.
+Figure 2 illustrates examples of the symmetry entropy, with the horizontal axes representing the values of $\Phi$ (left column) and $|\Phi|$ (right column), and the vertical axes corresponding to $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column).
+From the time-reversal antisymmetric condition [Eq.
+(7)], if $\Phi=a$ exists ($a>0$), then $\Phi=-a$ also exists, which is regarded as a pair of the observable.
+Suppose that the distribution $P(\Phi)$ is symmetric with respect to $\Phi=0$ for all of the pairs, as depicted in Fig.
+2(a).
+For this case, $P(|\Phi|)$ is very different from the original distribution $P(\Phi)$, which results in $\Lambda[P(\Phi)]=\ln 2$.
+In contrast, for an asymmetric distribution illustrated in Fig.
+2(b), $P(\Phi)$ and $P(|\Phi|)$ are the same, and thus $\Lambda[P(\Phi)]$ reduces to 00.
+Similarly, $\Lambda[P(\Phi)]$ is 00 for the bias pattern represented in Fig.
+2(c).
+When $\Phi(\Gamma)$ is a binary function (outputs two categories), we have $\Lambda[P(\Phi)]=H[P(\Phi)]$.
+For simplicity, let us consider the case where the observable $\Phi(\Gamma)$ does not take the value 00, a condition which holds in several problem settings.
+For instance, we may consider binary classification based on trajectories of stochastic processes.
 In this case, the observable $\Phi(\Gamma)$ does not include 00.
@@
 Later, we will consider the case where $\Phi(\Gamma)=0$ is included.
-Specifically, it can be proved that the following relation holds:
-$C[P(\Phi),P(-\Phi)]\geq\ln 2+H[P(|\Phi|)],$ (11)
-whose derivation is shown in Appendix A. In Eq. (11), $P(|\Phi|)$ is the probability distribution of $|\Phi|$, which is the absolute random variable of $\Phi$. Let $\Lambda[P(\Phi)]$ be
-$\Lambda[P(\Phi)]\equiv H[P(\Phi)]-H[P(|\Phi|)],$ (12)
-which is the entropy difference between the original distribution $P(\Phi)$ and its absolute valued distribution $P(|\Phi|)$.
-Here, we call $\Lambda[P(\Phi)]$ as symmetry entropy.
-The symmetry entropy $\Lambda[P(\Phi)]$ quantifies the extent of symmetry of $P(\Phi)$.
-Figure 1 depicts examples of values of the symmetry entropy, where the horizontal axes are values of $\Phi$ (left column) and $|\Phi|$ (right column) and the vertical axes are $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column).
-From the condition of time reversal [Eq.
-(9)], if $\Phi=a$ exists ($a>0$), then $\Phi=-a$ also exists, which is regarded as a pair of the observable.
-Suppose that the distribution $P(\Phi)$ is symmetric for all of the pairs, as depicted in Fig.
-1(a).
-For this case, $P(|\Phi|)$ is very different from the original distribution $P(\Phi)$, which results in $\Lambda[P(\Phi)]$ being $\ln 2$.
-In contrast, for an asymmetric distribution illustrated in Fig.
-1(b), $P(\Phi)$ and $P(|\Phi|)$ are the same, and thus $\Lambda[P(\Phi)]$ reduces to 00.
-To be more specific, the symmetry entropy quantifies how biased the probability distributions of pairs of observables are.
-When there is a strong bias in the probabilities of the pairs, the value of $\Lambda[P(\Phi)]$ becomes small.
-For example, $\Lambda[P(\Phi)]$ becomes 00 not only in cases where $P(\Phi)$ and $P(|\Phi|)$ are the same, as shown in Fig.
-1(b), but also in cases like Fig.
-1(c).
 It can be shown that
 $0\leq\Lambda[P(\Phi)]\leq\ln 2,$ (9)
-whose proof is provided in Appendix B.
+whose proof is provided in the End Matter.
 In Eq.
 (9), $\Lambda[P(\Phi)]$ being 00 and $\ln 2$ corresponds to asymmetric and symmetric distributions, respectively.
-Substituting Eq.
-(11) into Eq.
-(10), we obtain the trade-off between the entropy production and the asymmetry of $P(\Phi)$ quantified by entropy:
+Using the symmetry entropy $\Lambda[P(\Phi)]$, we obtain the trade-off between the entropy production and the asymmetry of $P(\Phi)$ quantified by entropy:
 $\Sigma\geq\ln 2-\Lambda[P(\Phi)]\geq 0.$ (10)
 Equation (10) is the main result of this study and referred to as thermodynamic entropic uncertainty relation.
+The derivation is shown in the End Matter.
 The right-hand side of Eq.
-(14) quantifies the asymmetry of probability distribution $P(\Phi)$.
-Therefore, Eq.
-(14) shows that, for arbitrary observable $\Phi$ satisfying the time reversal condition [Eq.
-(9)], the system requires the entropy production no less than $\ln 2-\Lambda[P(\Phi)]$.
-The trade-off between entropy production and observable asymmetry parallels traditional thermodynamic uncertainty relations, which illustrate a trade-off between the variance of observables and entropy production.
+(10) quantifies the asymmetry of the probability distribution $P(\Phi)$.
+Equation (10) demonstrates that for any observable $\Phi$ that satisfies the time-reversal antisymmetric condition specified in Eq.
+(7), the entropy production must be at least $\ln 2-\Lambda[P(\Phi)]$.
+The trade-off between entropy production and observable asymmetry parallels traditional thermodynamic uncertainty relations, which illustrate a trade-off between the relative variance of observables and entropy production.
 The thermodynamic uncertainty relations consider the relative variance $\mathrm{Var}[\Phi]/\mathbb{E}[\Phi]^{2}$.
@@
 (9), Eq.
-(14) can be regarded as a refinement of the second law using the entropy of the observable.
+(10) can be regarded as a refinement of the second law, employing the entropy of the observable.
 There are some advantages in employing the entropy instead of the variance.
-When dealing with the variance, the observable must yield a real number.
-However, in cases where the observable consists of classifications such as “success” and “failure”, the variance is not suitable.
+When the observable is a categorical variable that outputs labels such as “yes” or “no”, which is the case in classification tasks, the use of variance is not appropriate.
 Even in such cases, the Shannon entropy is well defined, indicating that Eq.
 (10) can be applied.
-Thus far, we assumed that $\Phi(\Gamma)$ takes values in a countable set, that is, $P(\Phi)$ is a discrete distribution.
-Consider the case where $\Phi(\Gamma)$ produces continuous values, where the summation should be replaced by the integration.
+The right-hand side of Eq.
+(10) can also be interpreted as a divergence.
+If we denote the Jensen-Shannon divergence by $\mathrm{JS}[P(X)\|P(Y)]$, then
+$\mathrm{JS}[P(\Phi)\|P(-\Phi)]=\ln 2-\Lambda[P(\Phi)],$ (11)
+the derivation of which is shown in the End Matter.
+Equation (10) provides a trade-off between the symmetry entropy and the entropy production.
+For a discrete probability distribution, $H[P(|\Phi|)]\geq 0$ holds.
+Therefore, $H[P(\Phi)]\geq\Lambda[P(\Phi)]$ and thus the following bound also holds:
+$\Sigma\geq\ln 2-H[P(\Phi)],$ (12)
+which purely relates the entropy production $\Sigma$ and the Shannon entropy of the observable $\Phi$.
+The right-hand side of Eq.
+(12) may not always be non-negative.
+Although the form of Eq.
+(12) is more appealing in terms of physical interpretation, the bound is weaker.
+When the observable $\Phi(\Gamma)$ is a binary function, the right-hand side of Eq.
+(12) becomes non-negative as $H[P(\Phi)]=\Lambda[P(\Phi)]$ for the binary case.
+So far, we assumed that $\Phi(\Gamma)$ takes values in a countable set, that is, $P(\Phi)$ is a discrete distribution.
+Consider the case where $\Phi(\Gamma)$ produces continuous values, where the summation should be replaced by an integration.
 When considering a continuous distribution, the notable difference is that the differential entropy may take negative values.
@@
 It is straightforward to extend the result to the case where the observable includes $\Phi(\Gamma)=0$.
-Specifically, when $\Phi(\Gamma)=0$ is included, the range of $\Lambda[P(\Phi)]$ is modified as follows:
+Specifically, when $\Phi(\Gamma)$ can equal 00, the range of $\Lambda[P(\Phi)]$ is modified as follows:
 $0\leq\Lambda[P(\Phi)]\leq[1-P(\Phi=0)]\ln 2$ (13)
@@
 (6), Eq.
-(16) holds for arbitrary observable $\Phi(\Gamma)$ satisfying Eq.
+(14) holds for any observable $\Phi(\Gamma)$ satisfying Eq.
 (7).
@@
 For such empty dynamics, the entropy production is 00, showing that both sides of Eq.
-(16) equal 00 and thus the inequality becomes equality.
-This consideration shows that $\Phi=0$ should be handled separately.
+(14) equal 00 and thus the inequality becomes an equality.
 However, note that this exceptional handling of $\Phi=0$ is not limited to Eq.
@@
 (1)], when the expectation of the current vanishes, the inequality is ill-defined.
-When $P(\Phi)$ is a probability density that is smooth around 00, the measure becomes $P(\Phi=0)=0$.
-Only when $P(\Phi)$ includes the delta-peaked contribution at $\Phi=0$, we use Eq.
-(16)
-Let us comment on the relation between Eq.
+Equations (10) and (14) saturate if and only if the system in equilibrium.
+For a detailed discussion on the saturation condition, see the supplementary material [26].
+Again, when $\Phi(\Gamma)$ is a discrete distribution, the following relation holds as well:
+$\Sigma\geq[1-P(\Phi=0)]\ln 2-H[P(\Phi)].$ (15)
+We remark on the relation between Eq.
 (14) and the Landauer principle [27].
-When considering a process that resets to one state from a state that exists with equal probability in two states, the following relation is obtained.
-$\Delta S_{m}\geq\ln 2,$ (17)
-where $\Delta S_{m}$ is the entropy increase in the surrounding medium.
-Note that the Landauer principle can be considered as a specific case of the second law of thermodynamics.
-Given the assumption that the system is in steady state, the change in system entropy $\Delta S$ is zero; consequently, $\Sigma=\Delta S_{m}$, where $\Delta S_{m}$ represents the increase in entropy in the environment.
+Consider a process that resets a system, which has an equal probability of being in one of two states, to a single specific state.
+The following relation is obtained:
+$\Delta S_{\mathrm{m}}\geq\ln 2,$ (16)
+where $\Delta S_{\mathrm{m}}$ is the entropy increase in the surrounding medium.
+Note that the Landauer principle can be considered as a specific case of the second law of thermodynamics [28].
+Given the assumption that the system is in steady state, the change in system entropy $\Delta S$ is zero; consequently, $\Sigma=\Delta S_{\mathrm{m}}$.
 Then, Eq.
-(14) can be expressed by $\Delta S_{m}\geq\ln 2-\Lambda[P(\Phi)]$.
+(10) can be expressed by $\Delta S_{\mathrm{m}}\geq\ln 2-\Lambda[P(\Phi)]$.
 When the probability distribution of $P(\Phi)$ ends up with a totally asymmetric distribution, i.e., $\Lambda[P(\Phi)]=0$, Eq.
@@
 (16) arises from a reduction in the Shannon entropy within the system’s state, implying that the Landauer principle is relevant when the system state changes over the time evolution.
-This contrast with Eq.
-(14), where the system is assumed to be steady state, and $\ln 2$ on its right-hand side arises from the asymmetry of the observable, not the system state.
-Example.—To illustrate Eq.
-(14), consider a one-dimensional biased random walk as shown in Fig.
-2(a).
-We consider a discrete time model, where the transition probabilities for moving right and left are denoted as $w_{+}$ and $w_{-}$, respectively, with the constraint $w_{+}+w_{-}=1$.
-The biased random walk plays an important role in biological modeling.
-For example, a biased random walk can model molecular motors, such as F1-ATPase [31] (Fig.
-2(b)).
-F1-ATPase is an enzyme that can rotate in discrete steps of $32\pi/$ 3 radians in the clockwise or counterclockwise direction [32].
-When rotating counterclockwise, F1-ATPase hydrolyzes ATP, while when rotating clockwise, F1-ATPase synthesizes ATP.
-The rotation in both directions is stochastic, which means that it occurs randomly.
-The counter-clockwise rotation can be considered as moving in the positive direction in the biased random walk, where the motor has a higher probability of rotating in the counter-clockwise direction.
-In the biased random walk, we consider the observable $\Phi(\Gamma)$, which takes either “$+$” or “--”.
-After the $L$ steps, the observable $\Phi$ is determined by comparing the final position to the initial position.
-If the final position is greater than the initial position, $\Phi$ is assigned the value “$+$”.
-Conversely, if the final position is less than or equal to the initial position, $\Phi$ is assigned the value “--”.
-When the number of steps is odd, there is no possibility of $\Phi(\Gamma)=0$.
-When considering the aforementioned molecular motor, this observable corresponds to checking whether the motor rotates in the counterclockwise direction or not.
-For details, see Appendix C.
-Figure 3: Results of the biased random walk.
-The sum of the entropy production and the symmetry entropy, $\Sigma+\Lambda[P(\Phi)]$, is plotted as a function of $w_{+}$ with the solid line.
-The dotted line denotes $\ln 2$, which is the lower bound of $\Sigma+\Lambda[P(\Phi)]$ in Eq.
-(14).
-Each of these terms, $\Sigma$ and $\Lambda[P(\Phi)]$, is plotted by the dashed line and the dot-dashed line, respectively.
-Here, we employ $L=5$.
-We calculate quantities, such as $\Sigma$ and $\Lambda[P(\Phi)]$, as a function of $w_{+}$ in Fig.
-3.
-Note that, for the binary case where $\Phi(\Gamma)$ takes two values, $\Lambda[P(\Phi)]=H[P(\Phi)]$.
+This contrasts with Eq.
+(10), where the system is assumed to be in steady state, and $\ln 2$ on its right-hand side arises from the asymmetry of the observable, not the state of the system.
+Example.—We demonstrate an application of the thermodynamic entropic uncertainty relation.
+Decision-making, a fundamental aspect of information processing, is crucial in diverse fields ranging from biology to information science.
+To illustrate this, we consider the diffusion decision model [29, 30, 31], which describes decision-making as a stochastic process of evidence accumulation.
+Let $\Gamma$ be a stochastic trajectory of the decision model within the time interval $[0,\tau]$ and $\Phi(\Gamma)$ be a function of $\Gamma$.
+In the diffusion decision model, a decision occurs when the accumulated evidence during $[0,\tau]$ surpasses thresholds.
+We assume that if the evidence is greater than or equal to $\theta$, then $\Phi=+1$ (‘yes”‘ class); if it is less than or equal to $-\theta$, then $\Phi=-1$ (“no” class); and in all other cases, $\Phi=0$ (neutral class).
+Figure 3(a) describes an example of the decision process.
+Here, $+1,-1,0$ are simply labels, so we can assign other labels.
+Specifically, let us consider the following Langevin dynamics for modeling the evidence accumulation often employed in the literature [29, 30, 31]:
+$\frac{dx}{dt}=f+\sqrt{2A}\xi(t),$ (17)
+where $f\in\mathbb{R}$ is a constant force, $\xi(t)$ is the white Gaussian noise with the correlation $\braket{\xi(t)\xi(t^{\prime})}=\delta(t-t^{\prime})$, and $A$ is the noise intensity.
+We define the accumulated evidence as $y(\tau)=\int_{0}^{\tau}\partial_{t}x(t)dt$ and the observable is $\Phi(\Gamma)=1$, $-1$, and 00 when $y(\tau)\geq\theta$, $y(\tau)\leq-\theta$, and $-\theta<y(\tau)<\theta$, respectively.
+Here, our interest is in the relationship between the entropy of the selected decision $H[P(\Phi)]$ and the entropy production of the Langevin equation.
+Since $\Phi(\Gamma)$ satisfies the time-reversal antisymmetric condition [Eq.
+(7)], it satisfies Eqs.
+(14) and (15), implying that greater entropy production is necessary for lower decision entropy $H[P(\Phi)]$ under the steady-state condition.
+In decision making, it is well known that there exists a speed-accuracy tradeoff [32, 33], responding faster often reduces accuracy, while achieving higher accuracy lengthens decision time.
+Through the thermodynamic entropic uncertainty relation, we show quantitatively that another tradeoff, an energy-accuracy tradeoff, holds in the decision model.
+We perform numerical calculations for the diffusion decision model [Eq.
+(17)] to confirm Eq.
+(15) [26].
+We calculate quantities, such as $\Sigma$ and $H[P(\Phi)]$, as a function of $f$ in Fig.
+3 for different threshold settings (a) $\theta=0$ and (b) $\theta=0.5$.
 We consider the steady-state condition and calculate the probability distribution $P(\Phi)$ and $\Sigma$.
 In Fig.
-3, $\Sigma+\Lambda[P(\Phi)]$, the sum of the entropy production $\Sigma$ and the symmetry entropy $\Lambda[P(\Phi)]$, is represented by the solid line, where the dotted line indicates $\ln 2$.
-As can be seen in the figure, $\Sigma+\Lambda[P(\Phi)]$ is always above $\ln 2$, confirming the validity of Eq.
+3, $\Sigma+H[P(\Phi)]$ is represented by the solid line, where the dotted line indicates (a) $\ln 2$ or (b) $(1-P(0))\ln 2$.
+As shown in Figs.
+3(a) and (b), $\Sigma+H[P(\Phi)]$ is always above $\ln 2$ or $(1-P(0))\ln 2$, respectively, confirming the validity of Eq.
 (15).
-To examine how each of these components, $\Sigma$ and $\Lambda[P(\Phi)]$, behave as a function of $w_{+}$, $\Sigma$ and $\Lambda[P(\Phi)]$ are represented by the dashed line and the dot-dashed line, respectively.
-Regarding entropy production $\Sigma$, it falls below $\ln 2$ in the region where $w_{+}$ ranges from $0.5$ to $0.65$.
-On the other hand, the symmetry entropy of the observable $\Lambda[P(\Phi)]$ is decreasing as a function of $w_{+}$, and except for $w_{+}=0.5$, it is below $\ln 2$.
-From these results, we can conclude that the inequality $\Sigma+\Lambda[P(\Phi)]\geq\ln 2$ holds only when considering the sum, $\Sigma+\Lambda[P(\Phi)]$.
-Conclusion.—In this study, we established a thermodynamic entropic uncertainty relation that links entropy production and the Shannon entropy of the observables.
+It is clear that in the case of $\theta=0$, the values of $\Sigma+H[P(\Phi)]$ and the lower bound are closer than in the case of $\theta=0.5$.
+To examine how each of these components, $\Sigma$ and $H[P(\Phi)]$, behave as a function of $f$, $\Sigma$ and $H[P(\Phi)]$ are represented by the dashed line and the dot-dashed line, respectively.
+Regarding entropy production $\Sigma$, it falls below $\ln 2$ in the region where $f$ ranges from $0.0$ to $0.25$ in Fig.
+3(a).
+On the other hand, the entropy of the observable $H[P(\Phi)]$ is decreasing as a function of $f$, and except for $f=0.0$, it is below $\ln 2$.
+From these results, we can conclude that the inequality $\Sigma+H[P(\Phi)]\geq\ln 2$ holds only when considering the sum, $\Sigma+H[P(\Phi)]$, that is, $\Sigma$ or $H[P(\Phi)]$ alone falls below $\ln 2$.
+Conclusion.—In this study, we established the thermodynamic entropic uncertainty relation that links the entropy production and the Shannon entropy of the observable.
 Our findings extend conventional thermodynamic uncertainty relations by incorporating measures based on entropy, highlighting the role of Shannon entropy of observables in stochastic thermodynamics.
 The derived inequality formalizes a fundamental trade-off between entropy production and the asymmetry of the observable distribution.
-We validated this relation using a biased random walk model.
+Through the application to the diffusion decision model, we revealed that accurate decision necessarily require higher entropy production.
 This framework provides a deeper understanding of nonequilibrium thermodynamics and expands the application of entropy-based uncertainty relations in stochastic systems.
-One direction of expansion is towards quantum systems.
-In recent years, the thermodynamic uncertainty relations in quantum systems [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 14, 43] have garnered significant attention.
+We have considered the Kullback-Leibler divergence with respect to trajectories $\Gamma$, but it is also possible to start from the divergence between different quantities (see the supplementary material [26]).
+Another direction of expansion is towards quantum systems.
+In recent years, the thermodynamic uncertainty relations in quantum systems [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47] have garnered significant attention.
 In particular, the thermodynamic uncertainty relations within the framework of continuous measurement are closely related to those in classical stochastic processes.
 This direction presents future challenges.
+Acknowledgements.
+This work was supported by JSPS KAKENHI Grant Number JP23K24915.
 Appendix A Derivation of the main result [Eq. (14)]
-A.1 Discrete case
-In this section, we derive Eq. (16). Let us introduce the absolute random variable of $\Phi$, which is expressed by $|\Phi|$. The probability distribution of $|\Phi|$ is given by
-$P(|\Phi|=\phi)=\begin{cases}P(\Phi=\phi)+P(\Phi=-\phi)&\phi\neq 0\\ P(\Phi=0)&\phi=0\end{cases}.$ (18)
+Let us introduce the absolute random variable of $\Phi$, which is expressed by $|\Phi|$. The probability distribution of $|\Phi|$ is given by
+$P(|\Phi|=\phi)=\begin{cases}P(\Phi=\phi)+P(\Phi=-\phi)&\phi>0\\ P(\Phi=0)&\phi=0\end{cases}$ (18)
 Therefore, the Shannon entropy of $|\Phi|$ is given by
@@
 $=-\sum_{\phi}P(\phi)\ln[P(\phi)+P(-\phi)]+P(0)\ln 2,$ (19)
-where we abbreviated $P(\phi)=P(\Phi=\phi)$. Then we compute $H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$ as follows:
+where we abbreviated $P(\phi)=P(\Phi=\phi)$.
+Using the monotonicity of Kullback-Leibler divergence [Eq. (5)] and the time-reversal property of $\Phi$ [Eq. (7)], we have
+$\Sigma$ $\geq D[P(\Phi)\|P(-\Phi)]$
+$=-H[P(\Phi)]+C[P(\Phi),P(-\Phi)].$ (20)
+Here, the cross entropy is evaluated as $C[P(\Phi),P(-\Phi)]=-\sum_{\phi}P(\Phi=\phi)\ln P(\Phi=-\phi)$.
+Since the cross entropy term is no less than the entropy, it is non-negative, $C[P(\Phi),P(-\Phi)]\geq H[P(\Phi)]\geq 0$.
+An important observation is that $C[P(\Phi),P(-\Phi)]$ is even bounded from below by a positive term.
+We calculate $H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$ as follows:
 $H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$
@@
 $\leq C[P(\Phi),P(-\Phi)].$ (23)
-Equation (22) proves Eq. (11) in the main text.
-A.2 Continuous case
-Next, we consider the continuous case. Basically, the derivation is the same as in the discrete case except that the summation should be replaced by the integration.
-Assume $P(\Phi)$ represents a probability density that is differentiable for all $\Phi$. For the continuous case, the probability density $P(|\Phi|)$ is defined by
+Substituting Eq. (23) into Eq. (20), Equation (23) proves Eqs. (10) and (14) in the main text.
+Next, we consider the continuous case.
+Basically, the derivation is the same as in the discrete case except that summation is replaced by integration.
+Assume $P(\Phi)$ represents a probability density that is differentiable for all $\Phi$.
+The Shannon entropy is defined by
+[…]
+For the continuous case, the probability density $P(|\Phi|)$ is defined by
 $P(|\Phi|=\phi)=P(\Phi=\phi)+P(\Phi=-\phi).$ (25)
@@
 When $P(\Phi)$ includes the contribution of the delta function in $\Phi=0$, this is not the case.
-The continuous entropy is given by
-$H[P(|\Phi|)]$
-[…]
-[…]
-[…]
-Again, we calculate $H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$ as follows:
-$H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$
-[…]
-[…]
-Using Eq. (25), we prove
+Following the same procedure as the discrete case, we obtain
 $H[P(|\Phi|)]+\ln 2$ $\leq C[P(\Phi),P(-\Phi)].$ (26)
-Appendix B Proof of Eq. (13)
+Appendix B Symmetry entropy and Jensen-Shannon divergence
+The Jensen-Shannon divergence is defined by
+$\mathrm{JS}[P(X)\|P(Y)]$ $\equiv\frac{1}{2}D\left[P(X)\|\frac{P(X)+P(Y)}{2}\right]$
+$+\frac{1}{2}D\left[P(Y)\|\frac{P(X)+P(Y)}{2}\right],$ (27)
+which satisfies
+$0\leq\mathrm{JS}[P(X)\|P(Y)]\leq\ln 2$ (28)
+Using the Jensen-Shannon divergence, we have
+$\mathrm{JS}[P(\Phi)\|P(-\Phi)]$
+$=-\sum_{\phi}\frac{P(\phi)+P(-\phi)}{2}\ln\frac{P(\phi)+P(-\phi)}{2}$
+$+\frac{1}{2}\sum_{\phi}P(\phi)\ln P(\phi)+\frac{1}{2}\sum_{\phi}P(-\phi)\ln P(-\phi)$
+$=\ln 2-\sum_{\phi}P(\phi)\ln\left[P(\phi)+P(-\phi)\right]+\sum_{\phi}P(\phi)\ln P(\phi)$
+$=\ln 2+H[P(|\Phi|)]-P(0)\ln 2-H[P(\Phi)]$
+$=\ln 2-P(0)\ln 2-\Lambda[P(\Phi)],$ (29)
+where we used Eq. (19). Equation (29) is Eq. (11) in the main text.
+Appendix C Proof of Eq. (9)
 In this section, we prove Eq. (9). Here, we show the relation for $P(\Phi=0)\geq 0$. The relation which we want to show is given by
@@
 $H[P(X),P(Y)]$ $=H[P(X)|P(Y)]+H[P(Y)]$
-$=H[P(Y)|P(X)]+H[P(X)].$ (28)
+$=H[P(Y)|P(X)]+H[P(X)],$ (31)
+where $H[P(Y)|P(X)]$ is the conditional entropy:
+$H[P(Y)|P(X)]$
+$\equiv-\sum_{x,y}P(X=x,Y=y)\ln P(Y=y|X=x).$ (32)
 Substituting $X=\Phi$ and $Y=|\Phi|$, we obtain
@@
 However, we can show that the first inequality part also holds for the continuous case by a direct calculation.
+Specifically, $H[P(\Phi)]-H[P(|\Phi|)]$ is
+$H[P(\Phi)]-H[P(|\Phi|)]$
+[…]
+[…]
+$=\int_{0}^{\infty}d\phi\,\Bigl{[}[P(\phi)+P(-\phi)]\ln\left[P(\phi)+P(-\phi)\right]$
+$-P(\phi)\ln P(\phi)-P(-\phi)\ln P(-\phi)\Bigr{]}.$ (34)
+Because $(a+b)\ln(a+b)-a\ln a-b\ln b>0$ for $a>0$ and $b>0$, $H[P(\Phi)]-H[P(|\Phi|)]\geq 0$ is proved.
 Next, we prove the second part of the inequality of Eq. (30), which can be done following the same approach as in Appendix A. Using Eq. (19), we have
@@
 $\geq-\ln 2+P(0)\ln 2,$ (35)
-where we again used the Jensen inequality.
-Equation (30) proves the second inequality part of Eq.
+where we again used the Jensen inequality. Equation (35) proves the second inequality part of Eq. (30).
+By adopting the formulation based on the Jensen-Shannon divergence [Eq.
+(11)], the second part of the inequality in Eq.
+(30) can be directly derived from Eq.
 (28).
-For the continuous case, by following the same argument by replacing the summation with the integration, the following relation holds:
-$H[P(|\Phi|)]-H[P(\Phi)]$ $\geq-\ln 2.$ (31)
-Appendix C Details of example
-We show details of the biased random walk example. The probability distribution at position $x$ after $L$ steps is given by
-$p(x)=\begin{cases}{\displaystyle\binom{L}{\frac{L+x}{2}}p^{\frac{L+x}{2}}(1-p)^{\frac{L-x}{2}}}&L+x\in\mathrm{even}\\ 0&\mathrm{otherwise}\end{cases}.$ (32)
-Using $p(x)$ in Eq.
-(32), for odd $L$, the probability of $\Phi$ is calculated by $P(\Phi=\mathrm{decrease})=\sum_{x=-L}^{-1}p(x)$ and $P(\Phi=\mathrm{increase})=\sum_{x=1}^{L}p(x)$, .
-which can be used to calculate the Shannon entropy of $P(\Phi)$.
-Moreover, under the steady-state condition, the entropy production $\Sigma$ up to $L$ steps is given by
-$\Sigma=L(w_{+}-w_{-})\ln\frac{w_{+}}{w_{-}}.$ (33)
+However, note that the first inequality part of Eq.
+(30) cannot be obtained from Eq.
+(28).

### v5

@@
 As an application, we demonstrate our relation in the diffusion decision model, revealing a fundamental trade-off between decision accuracy and entropy production in stochastic decision-making processes.
-Introduction.— The thermodynamic uncertainty relation [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] expresses a universal trade-off between the entropy production and the relative variance of time-integrated observables for systems operating far from equilibrium.
+I Introduction
+The thermodynamic uncertainty relation [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] expresses a universal trade-off between the entropy production and the relative variance of time-integrated observables for systems operating far from equilibrium.
 For a steady-state stochastic thermodynamic system and a trajectory observable $F$, the thermodynamic uncertainty relation states [1, 2]
@@
 where $\mathbb{E}[F]$ and $\mathrm{Var}[F]$ are the expectation value and variance of $F$, respectively, and $\Sigma$ denotes the entropy production during the observation interval.
-The thermodynamic uncertainty relation implies that achieving higher precision inevitably requires greater entropy production–a manifestation of the “no free lunch” principle in thermodynamics.
+The thermodynamic uncertainty relation implies that achieving higher precision inevitably requires greater entropy production, a manifestation of the “no free lunch” principle in thermodynamics.
 The thermodynamic uncertainty relation given by Eq.
@@
 However, even in such cases, we can still calculate the Shannon entropy.
-Therefore, to analyze the tradeoff between the accuracy of the categorical outputs and the entropy production, a thermodynamic uncertainty relation based on Shannon entropy is required.
+Therefore, to analyze the trade-off between the accuracy of the categorical outputs and the entropy production, a thermodynamic uncertainty relation based on Shannon entropy is required.
 The issue of how to quantify uncertainty has been actively studied in quantum mechanics.
@@
 Later, Hirschman showed that the variance can be replaced by the Shannon entropy [13, 14].
-Maassen and Uffink [15] derived a precise statement of this entropic uncertainty relation [16]: the sum of the entropies of the two observables cannot fall below a bound determined by the extent of their eigenstate overlap.
+Maassen and Uffink [15] derived a precise statement of this entropic uncertainty relation [16]; the sum of the entropies of the two observables cannot fall below a bound determined by the extent of their eigenstate overlap.
 The entropic uncertainty relation plays a fundamental role in quantum cryptography, especially in quantum key distribution protocols [17].
@@
 The trade-off relation in thermodynamic cost has been actively studied, particularly in terms of thermodynamic uncertainty relations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] using the variance of observable quantities and speed limits using the distance between states [18, 19, 20, 9, 21, 22, 23].
-This study demonstrates a trade-off between the entropy of the observable and the entropy production, which is expected to lead to the derivation of other trade-off relations.
+This study demonstrates the trade-off between the entropy of the observable and the entropy production, which is expected to lead to the derivation of other trade-off relations.
 Figure 1: Conceptual representation of the thermodynamic entropic uncertainty relation.
@@
 (b) Trajectory of the stochastic process shown in (a).
-$\Gamma$ denotes time evolution of a realization of the process.
+$\Gamma$ denotes the time evolution of a realization of the process.
 $\Gamma^{\dagger}$ is the time-reversal of $\Gamma$.
@@
 Horizontal axes denote values of $\Phi$ (left column) and $|\Phi|$ (right column).
-The values of $\Phi$ always form pairs; $-1$ and $1$, and $-2$ and $2$ are two pairs in the examples.
-Vertical axes denote probability distribution $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column).
-(a) Distribution symmetric with respect to $\Phi=0$ where probabilities of all of the pairs are identical.
+The values of $\Phi$ always form pairs; 1-1 and $1$, and 2-2 and $2$ are two pairs in the examples.
+Vertical axes denote the probability distribution $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column).
+(a) Distribution symmetric with respect to $\Phi=0$, where the probabilities of all of the pairs are identical.
 $P(|\Phi|)$ is different from the original distribution $P(\Phi)$ and thus $\Lambda[P(\Phi)]$ is $\ln 2$.
-(b) Asymmetric distribution where probabilities of all of the pairs are fully biased.
-$P(|\Phi|)$ and $P(\Phi)$ are identical and thus $\Lambda[P(\Phi)]$ is 00.
+(b) Asymmetric distribution where the probabilities of all of the pairs are fully biased.
+$P(|\Phi|)$ and $P(\Phi)$ are identical and thus $\Lambda[P(\Phi)]$ is 0.
 (c) Asymmetric distribution where probabilities of all of the pairs are fully biased.
-The distributions $P(|\Phi|)$ and $P(\Phi)$ are different, but they effectively become the same if the labels for the pair $-1$ and $1$ are swapped.
+The distributions $P(|\Phi|)$ and $P(\Phi)$ are different, but they effectively become the same if the labels for the pair 1-1 and $1$ are swapped.
 Therefore, they are essentially the same distribution.
 This results in $\Lambda[P(\Phi)]=0$.
-Methods.—Let $X$ be a random variable and $P(X)$ be its probability distribution. The Shannon entropy $H[P(X)]$ is defined by
+II Methods
+Let $X$ be a random variable and $P(X)$ be its probability distribution. The Shannon entropy $H[P(X)]$ is defined by
 $H[P(X)]\equiv-\sum_{x}P(X=x)\ln P(X=x).$ (2)
@@
 $D[P(X)\|P(Y)]$ $\equiv\sum_{x}P(X=x)\ln\frac{P(X=x)}{P(Y=x)}$
-$=-H[P(X)]+C[P(X),P(Y)].$ (3)
+$=-H[P(X)]+C[P(X),P(Y)],$ (3)
 where $C[P(X),P(Y)]$ is the cross entropy:
@@
 Then the following monotonicity relation holds:
-$D[P(X)\|P(Y)]\geq D[P(\tilde{X})\|P(\tilde{Y})],$ (5)
-where $P(\tilde{X})$ and $P(\tilde{Y})$ are probability distributions of the transformed variables $\tilde{X}$ and $\tilde{Y}$, respectively.
+[…]
+where P(X~)P($\tilde${X}) and P(Y~)P($\tilde${Y}) are probability distributions of the transformed variables $\tilde{X}$ and $\tilde{Y}$, respectively.
 Having introduced basic concepts of the divergence, we move to consider stochastic thermodynamic systems.
-Stochastic thermodynamics considers processes whose state change is described by a stochastic process (Fig.
+Stochastic thermodynamics considers processes whose state changes are described by a stochastic process (Fig.
 1(a)).
@@
 $\Sigma=D[\mathcal{P}(\Gamma)\|\mathcal{P}(\Gamma^{\dagger})].$ (6)
-Note that the expression of Eq.
+Note that the expression in Eq.
 (6) also holds for continuous processes such as Langevin dynamics.
@@
 However, most of the results below hold for the continuous case as well.
-Figure 3: (a) Description of the diffusion decision model.
-The green trajectory illustrates a realization that exceeds the threshold $\theta$, resulting in a decision of $+1$ (“yes” class).
-Conversely, the purple trajectory represents a realization that falls below $-\theta$, leading to a decision of $-1$ (“no” class).
-(b) and (c) Entropies as a function of parameter $f$.
-In both figures, the blue dashed line shows the entropy production $\Sigma$, and the purple dot-dashed line represents $H[P(\Phi)]$, the entropy of the distribution $P(\Phi)$.
-Their sum, $\Sigma+H[P(\Phi)]$, is also depicted by the red solid line.
-The orange dotted line represents $\ln 2$ and $(1-P(0))\ln 2$ in (a) and (b), respectively, which serves as the lower bound for $\Sigma+H[P(\Phi)]$ according to Eq.
-(15).
-These plots are shown for $\theta=0$ in (b) and $\theta=0.5$ in (c), with other parameters being $A=1$ and $\tau=10$.
-Results.—We derive the thermodynamic entropic uncertainty relation, which is the main result of this Letter.
-We first introduce an entropic measure which quantifies the observable of a trajectory $\Gamma$.
-Let $\Lambda[P(\Phi)]$ be
+III Results
+We first introduce an entropic measure that quantifies a trajectory observable. Let $\Lambda[P(\Phi)]$ be
 $\Lambda[P(\Phi)]\equiv H[P(\Phi)]-H[P(|\Phi|)],$ (8)
-which is the entropy difference between the original distribution $P(\Phi)$ and its absolute valued distribution $P(|\Phi|)$.
+which is the entropy difference between the original distribution $P(\Phi)$ and its absolute-value distribution $P(|\Phi|)$.
 Here, we call $\Lambda[P(\Phi)]$ symmetry entropy.
@@
 Intuitively speaking, the difference $H[P(\Phi)]-H[P(|\Phi|)]$ quantifies how much “sign” information remains in the distribution $P(\Phi)$ once its magnitude $P(|\Phi|)$ is known.
-If $P(\Phi)$ is perfectly symmetric with respect to $\Phi=0$, knowing $P(|\Phi|)$ tells us nothing about the sign, so $\Lambda[P(\Phi)]$ reaches its maximum of $\ln 2$ (one bit).
-In contrast, if $P(\Phi)$ is fully biased in one direction, the sign is already determined by $P(|\Phi|)$ and $\Lambda[P(\Phi)]$ vanishes.
-Figure 2 illustrates examples of the symmetry entropy, with the horizontal axes representing the values of $\Phi$ (left column) and $|\Phi|$ (right column), and the vertical axes corresponding to $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column).
-From the time-reversal antisymmetric condition [Eq.
-(7)], if $\Phi=a$ exists ($a>0$), then $\Phi=-a$ also exists, which is regarded as a pair of the observable.
-Suppose that the distribution $P(\Phi)$ is symmetric with respect to $\Phi=0$ for all of the pairs, as depicted in Fig.
-2(a).
-For this case, $P(|\Phi|)$ is very different from the original distribution $P(\Phi)$, which results in $\Lambda[P(\Phi)]=\ln 2$.
-In contrast, for an asymmetric distribution illustrated in Fig.
-2(b), $P(\Phi)$ and $P(|\Phi|)$ are the same, and thus $\Lambda[P(\Phi)]$ reduces to 00.
-Similarly, $\Lambda[P(\Phi)]$ is 00 for the bias pattern represented in Fig.
-2(c).
-When $\Phi(\Gamma)$ is a binary function (outputs two categories), we have $\Lambda[P(\Phi)]=H[P(\Phi)]$.
-For simplicity, let us consider the case where the observable $\Phi(\Gamma)$ does not take the value 00, a condition which holds in several problem settings.
+For simplicity, let us consider the case where the observable $\Phi(\Gamma)$ does not take the value 0, a condition which holds in several problem settings.
 For instance, we may consider binary classification based on trajectories of stochastic processes.
-In this case, the observable $\Phi(\Gamma)$ does not include 00.
+In this case, the observable $\Phi(\Gamma)$ does not include 0.
 $\Phi(\Gamma)=0$ should be handled separately because $\Phi(\Gamma)=\Phi(\Gamma^{\dagger})=0$ from Eq.
@@
 $0\leq\Lambda[P(\Phi)]\leq\ln 2,$ (9)
-whose proof is provided in the End Matter.
-In Eq.
-(9), $\Lambda[P(\Phi)]$ being 00 and $\ln 2$ corresponds to asymmetric and symmetric distributions, respectively.
+whose proof is provided in Appendix B.
+Let us elaborate on the necessary and sufficient conditions for achieving 0 and $\ln 2$ in Eq.
+(9).
+Figure 2 illustrates examples of the symmetry entropy, with the horizontal axes representing the values of $\Phi$ (left column) and $|\Phi|$ (right column), and the vertical axes corresponding to $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column).
+From the time-reversal antisymmetric condition [Eq.
+(7)], if $\Phi=\phi$ exists ($\phi>0$), then $\Phi=-\phi$ also exists, which is regarded as a pair of the observable.
+From the calculation in Appendix B, $\Lambda[P(\Phi)]=\ln 2$ if and only if $P(\Phi=\phi)=P(\Phi=-\phi)$ for all $\phi$.
+This corresponds to the symmetric probability distribution as shown in Fig.
+2(a).
+In contrast, $\Lambda[P(\Phi)]=0$ is satisfied if and only if $P(\Phi=\phi)=0$ or $P(\Phi=-\phi)=0$ for all $\phi$.
+This case corresponds to asymmetric probability distributions as shown in Figs.
+2(b) and (c).
+When $\Phi(\Gamma)$ is a binary function (outputs two categories), we have $\Lambda[P(\Phi)]=H[P(\Phi)]$.
 Using the symmetry entropy $\Lambda[P(\Phi)]$, we obtain the trade-off between the entropy production and the asymmetry of $P(\Phi)$ quantified by entropy:
 $\Sigma\geq\ln 2-\Lambda[P(\Phi)]\geq 0.$ (10)
-Equation (10) is the main result of this study and referred to as thermodynamic entropic uncertainty relation.
-The derivation is shown in the End Matter.
+Equation (10) is the main result of this study and is referred to as the thermodynamic entropic uncertainty relation.
+The derivation is shown in Appendix A.
 The right-hand side of Eq.
@@
 When the observable is a categorical variable that outputs labels such as “yes” or “no”, which is the case in classification tasks, the use of variance is not appropriate.
-Even in such cases, the Shannon entropy is well defined, indicating that Eq.
-(10) can be applied.
+Even in such cases, the symmetry entropy defined by Eq.
+(8) can handle categorical outputs.
+As mentioned, the outputs of the observables constitute pairs, i.e., $\Phi=\phi$ and $\Phi=-\phi$.
+As the value of $\phi$ does not affect the calculation of the entropy $H[P(\Phi)]$, the paired outputs can be replaced with categorical outputs such as $\Phi=\phi=\text{``yes''}$ or $\Phi=-\phi=\text{``no''}$.
 The right-hand side of Eq.
@@
 $\mathrm{JS}[P(\Phi)\|P(-\Phi)]=\ln 2-\Lambda[P(\Phi)],$ (11)
-the derivation of which is shown in the End Matter.
+the derivation of which is shown in Appendix C.
 Equation (10) provides a trade-off between the symmetry entropy and the entropy production.
@@
 Therefore, the main result of Eq.
-(10) is well defined for the continuous case as well.
-The derivation of Eq.
-(10) assumed that the observable $\Phi(\Gamma)$ does not include $\Phi(\Gamma)=0$.
+(10) holds for the continuous case as well.
+So far, we assumed that the observable $\Phi(\Gamma)$ does not include $\Phi(\Gamma)=0$.
 It is straightforward to extend the result to the case where the observable includes $\Phi(\Gamma)=0$.
-Specifically, when $\Phi(\Gamma)$ can equal 00, the range of $\Lambda[P(\Phi)]$ is modified as follows:
-$0\leq\Lambda[P(\Phi)]\leq[1-P(\Phi=0)]\ln 2$ (13)
+Specifically, when $\Phi(\Gamma)$ can equal 0, the range of $\Lambda[P(\Phi)]$ is modified as follows (see Appendix B):
+$0\leq\Lambda[P(\Phi)]\leq[1-P(\Phi=0)]\ln 2.$ (13)
 The bound becomes
@@
 (10) as the specific case $P(\Phi=0)=0$.
+The derivation is shown in Appendix A.
 As long as the entropy production is given by Eq.
@@
 (14), $\Phi=0$ plays a special role.
-When considering the process of doing nothing, the observable $\Phi$ is always 00 implying $P(\Phi=0)=1$.
-For such empty dynamics, the entropy production is 00, showing that both sides of Eq.
-(14) equal 00 and thus the inequality becomes an equality.
-However, note that this exceptional handling of $\Phi=0$ is not limited to Eq.
-(14).
-In the conventional thermodynamic uncertainty relation [Eq.
-(1)], when the expectation of the current vanishes, the inequality is ill-defined.
-Equations (10) and (14) saturate if and only if the system in equilibrium.
-For a detailed discussion on the saturation condition, see the supplementary material [26].
+When considering the process of doing nothing, the observable $\Phi$ is always 0, implying $P(\Phi=0)=1$.
+For such empty dynamics, the entropy production is 0, showing that both sides of Eq.
+(14) equal 0 and thus the inequality becomes an equality.
+Although the conventional thermodynamic uncertainty relation [Eq.
+(1)] does not handle $\Phi=0$ as an exceptional case, it becomes ill-defined for the equilibrium condition where both the expectation of the current and the entropy production vanish.
+Equations (10) and (14) saturate if and only if the system is in equilibrium.
+For a detailed discussion on the saturation condition, see Appendix D.
 Again, when $\Phi(\Gamma)$ is a discrete distribution, the following relation holds as well:
@@
 (10) can be expressed by $\Delta S_{\mathrm{m}}\geq\ln 2-\Lambda[P(\Phi)]$.
-When the probability distribution of $P(\Phi)$ ends up with a totally asymmetric distribution, i.e., $\Lambda[P(\Phi)]=0$, Eq.
+When the probability distribution $P(\Phi)$ is totally asymmetric, i.e., $\Lambda[P(\Phi)]=0$, Eq.
 (10) is formally identical to the Landauer principle.
-However, note that the right side of Eq.
+However, note that the right-hand side of Eq.
 (16) arises from a reduction in the Shannon entropy within the system’s state, implying that the Landauer principle is relevant when the system state changes over the time evolution.
@@
 (10), where the system is assumed to be in steady state, and $\ln 2$ on its right-hand side arises from the asymmetry of the observable, not the state of the system.
-Example.—We demonstrate an application of the thermodynamic entropic uncertainty relation.
+IV Example
+Figure 3: Description of the diffusion decision model.
+The green trajectory illustrates a realization that exceeds the threshold $\theta$, resulting in a decision of +1+1 (“yes” class).
+Conversely, the purple trajectory represents a realization that falls below $\theta$, leading to a decision of 1-1 (“no” class).
+Trajectories that do not fall into either category correspond to a decision of 0 (“neutral” class).
+Figure 4: Results for the diffusion decision model with the discrete-output observable [Eq.
+(18)].
+Panels (a) and (b) show entropic quantities as a function of parameter $f$ for $\theta=0$ $\theta=0.5$, respectively, with $A=1$ and $\tau=10$.
+Panel (c) shows entropic quantities as a function of $\tau$ with $\theta=0$, $A=1$, and $f=0.5$.
+In these figures, the blue dashed line shows the entropy production $\Sigma$, and the purple dot-dashed line represents $H[P(\Phi)]$.
+Their sum, $\Sigma+H[P(\Phi)]$, is also depicted by the red solid line.
+The orange dotted line represents $\ln 2$ in (a) and (c) and $(1-P(0))\ln 2$ in (b), which serves as the lower bound for $\Sigma+H[P(\Phi)]$ according to Eq.
+(15).
+We demonstrate an application of the thermodynamic entropic uncertainty relation.
 Decision-making, a fundamental aspect of information processing, is crucial in diverse fields ranging from biology to information science.
 To illustrate this, we consider the diffusion decision model [30, 31, 32], which describes decision-making as a stochastic process of evidence accumulation.
-Let $\Gamma$ be a stochastic trajectory of the decision model within the time interval $[0,\tau]$ and $\Phi(\Gamma)$ be a function of $\Gamma$.
-In the diffusion decision model, a decision occurs when the accumulated evidence during $[0,\tau]$ surpasses thresholds.
-We assume that if the evidence is greater than or equal to $\theta$, then $\Phi=+1$ (‘yes”‘ class); if it is less than or equal to $-\theta$, then $\Phi=-1$ (“no” class); and in all other cases, $\Phi=0$ (neutral class).
-Figure 3(a) describes an example of the decision process.
-Here, $+1,-1,0$ are simply labels, so we can assign other labels.
+Let $\Gamma$ be a stochastic trajectory of the decision model within the time interval $0,\tau]$ and $\Phi(\Gamma)$ be a function of $\Gamma$.
+In the diffusion decision model, a decision occurs when the accumulated evidence during $0,\tau]$ surpasses thresholds.
+We assume that if the evidence is greater than or equal to $\theta$, then $\Phi=+1$ (“yes” class); if it is less than or equal to $\theta$, then $\Phi=-1$ (“no” class); and in all other cases, $\Phi=0$ (neutral class).
+Figure 3 describes an example of the decision process.
+Here, +1,1,0+1,-1,0 are simply labels, so we can assign other labels.
 Specifically, let us consider the following Langevin dynamics for modeling the evidence accumulation often employed in the literature [30, 31, 32]:
@@
 where $f\in\mathbb{R}$ is a constant force, $\xi(t)$ is the white Gaussian noise with the correlation $\braket{\xi(t)\xi(t^{\prime})}=\delta(t-t^{\prime})$, and $A$ is the noise intensity.
-We define the accumulated evidence as $y(\tau)=\int_{0}^{\tau}\partial_{t}x(t)dt$ and the observable is $\Phi(\Gamma)=1$, $-1$, and 00 when $y(\tau)\geq\theta$, $y(\tau)\leq-\theta$, and $-\theta<y(\tau)<\theta$, respectively.
-Here, our interest is in the relationship between the entropy of the selected decision $H[P(\Phi)]$ and the entropy production of the Langevin equation.
+We define the accumulated evidence as $y(\tau)=\int_{0}^{\tau}\partial_{t}x(t)dt$ and the observable is
+$\Phi(\Gamma)=\begin{cases}+1\;\;(\text{yes})&{\displaystyle y(\tau)\geq\theta}\\ -1\;\;(\text{no})&{\displaystyle y(\tau)\leq-\theta}\\ 0\;\;(\text{neutral})&\mathrm{otherwise}\end{cases}.$ (18)
+Explicit expressions of the probability distribution of $y$ are shown in Appendix E. The entropy production during $0,\tau]$ is given by
+$\Sigma=\frac{\tau f^{2}}{A}.$ (19)
+The thermodynamic entropic uncertainty relation loses its predictive power when the entropy production $\Sigma$ exceeds the threshold $\ln 2$.
+For the diffusion decision model, this occurs when the time parameter $\tau$ exceeds the critical value $\tau^{*}$ defined as
+$\tau^{*}\equiv\frac{A}{f^{2}}\ln 2.$ (20)
+Here, our interest is in the relationship between the entropy of the selected decision $H[P(\Phi)]$ and the entropy production $\Sigma$ of the Langevin equation.
 Since $\Phi(\Gamma)$ satisfies the time-reversal antisymmetric condition [Eq.
@@
 (14) and (15), implying that greater entropy production is necessary for lower decision entropy $H[P(\Phi)]$ under the steady-state condition.
-In decision making, it is well known that there exists a speed-accuracy tradeoff [32, 33], responding faster often reduces accuracy, while achieving higher accuracy lengthens decision time.
-Through the thermodynamic entropic uncertainty relation, we show quantitatively that another tradeoff, an energy-accuracy tradeoff, holds in the decision model.
+In decision making, it is well known that there exists a speed-accuracy trade-off [33, 34], responding faster often reduces accuracy, while achieving higher accuracy lengthens decision time.
+Through the thermodynamic entropic uncertainty relation, we show quantitatively that another trade-off, an energy-accuracy trade-off, holds in the decision model.
 We perform numerical calculations for the diffusion decision model [Eq.
@@
 (15) [26].
-We calculate quantities, such as $\Sigma$ and $H[P(\Phi)]$, as a function of $f$ in Fig.
-3 for different threshold settings (a) $\theta=0$ and (b) $\theta=0.5$.
-We consider the steady-state condition and calculate the probability distribution $P(\Phi)$ and $\Sigma$.
-In Fig.
-3, $\Sigma+H[P(\Phi)]$ is represented by the solid line, where the dotted line indicates (a) $\ln 2$ or (b) $(1-P(0))\ln 2$.
+We calculate quantities, such as $\Sigma$ and $H[P(\Phi)]$, as a function of $f$ in Figs.
+4(a) and (b) for different threshold settings (a) $\theta=0$ and (b) $\theta=0.5$.
+We consider the steady-state condition.
+In Figs.
+4(a) and (b), $\Sigma+H[P(\Phi)]$ is represented by the solid line, where the dotted line indicates (a) $\ln 2$ or (b) $(1-P(0))\ln 2$.
 As shown in Figs.
-3(a) and (b), $\Sigma+H[P(\Phi)]$ is always above $\ln 2$ or $(1-P(0))\ln 2$, respectively, confirming the validity of Eq.
+4(a) and (b), $\Sigma+H[P(\Phi)]$ is always above $\ln 2$ or $(1-P(0))\ln 2$, respectively, confirming the validity of Eq.
 (15).
@@
 Regarding entropy production $\Sigma$, it falls below $\ln 2$ in the region where $f$ ranges from $0.0$ to $0.25$ in Fig.
-3(a).
+4(a).
 On the other hand, the entropy of the observable $H[P(\Phi)]$ is decreasing as a function of $f$, and except for $f=0.0$, it is below $\ln 2$.
 From these results, we can conclude that the inequality $\Sigma+H[P(\Phi)]\geq\ln 2$ holds only when considering the sum, $\Sigma+H[P(\Phi)]$, that is, $\Sigma$ or $H[P(\Phi)]$ alone falls below $\ln 2$.
-Conclusion.—In this study, we established the thermodynamic entropic uncertainty relation that links the entropy production and the Shannon entropy of the observable.
+We study the $\tau$ dependence of the entropic quantities.
+We calculate the entropic quantities, such as $\Sigma$ and $H[P(\Phi)]$, as a function of $\tau$ in Fig.
+4(c).
+The meaning of each line is the same as in Fig.
+4(a).
+Since the entropy production increases linearly as a function of $\tau$ [Eq.
+(19)], we see that the predictive power of the thermodynamic entropic uncertainty relation is lost around $\tau^{*}=2.8$ [Eq.
+(20)] for this parameter setting.
+Figure 5: Results for the diffusion decision model with the continuous-output observable [Eq.
+(21)].
+Panels (a) and (b) show the entropic quantities versus $f$ for $A=1$ and $A=0.1$, respectively, with $\tau$ fixed at $10$.
+The blue dashed line shows the entropy production $\Sigma$, and the purple dot-dashed line represents $\Lambda[P(\Phi)]$, the symmetry entropy.
+Their sum, $\Sigma+\Lambda[P(\Phi)]$, is also depicted with the red solid line.
+The orange dotted line displays $\ln 2$, which serves as the lower bound for $\Sigma+\Lambda[P(\Phi)]$ according to Eq.
+(14).
+Next, we consider the continuous-output observable case, where the observable is
+$\Phi(\Gamma)=y(\tau).$ (21)
+Equation (21) obviously satisfies the time-reversal antisymmetry condition.
+We numerically calculate the entropic quantities.
+We show $\Sigma$ and $\Lambda[P(\Phi)]$ as a function of $f$ in Fig.
+5 for different noise intensity settings (a) $A=1$ and (b) $A=0.1$.
+Figure 5 shows the entropic quantities: $\Sigma+\Lambda[P(\Phi)]$ (solid line), $\ln 2$ (dotted line), $\Lambda[P(\Phi)]$ (dot-dashed line), and $\Sigma$ (dashed line).
+Again, Figs.
+5(a) and (b) demonstrate that the sum $\Sigma+\Lambda[P(\Phi)]$ consistently exceeds $\ln 2$, supporting the correctness of Eq.
+(14).
+Specifically, when the noise intensity $A$ is set to the smaller value, as shown in Fig.
+5(b), the symmetry entropy $\Lambda[P(\Phi)]$ approaches zero as $f$ increases.
+When $A$ is smaller, the entropy production $\Sigma$ increases, leading to a rapid increase in the value of $\Sigma+\Lambda[P(\Phi)]$ as a function of $f$.
+In Fig.
+5(a), in the range where $f$ varies from $0.0$ to $0.25$, the entropy production $\Sigma$ is less than $\ln 2$.
+The symmetry entropy $\Lambda[P(\Phi)]$ decreases with increasing $f$ and remains below $\ln 2$ for all values except when $f=0.0$.
+Based on these observations, it is evident that the inequality $\Sigma+\Lambda[P(\Phi)]\geq\ln 2$ is satisfied only when considering the combined terms, $\Sigma+\Lambda[P(\Phi)]$.
+Individually, both $\Sigma$ and $\Lambda[P(\Phi)]$ drop below $\ln 2$.
+V Conclusion
+In this study, we established the thermodynamic entropic uncertainty relation that links the entropy production and the Shannon entropy of the observable.
 Our findings extend conventional thermodynamic uncertainty relations by incorporating measures based on entropy, highlighting the role of Shannon entropy of observables in stochastic thermodynamics.
 The derived inequality formalizes a fundamental trade-off between entropy production and the asymmetry of the observable distribution.
-Through the application to the diffusion decision model, we revealed that accurate decision necessarily require higher entropy production.
+Through the application to the diffusion decision model, we revealed that accurate decisions necessarily require higher entropy production.
 This framework provides a deeper understanding of nonequilibrium thermodynamics and expands the application of entropy-based uncertainty relations in stochastic systems.
-We have considered the Kullback-Leibler divergence with respect to trajectories $\Gamma$, but it is also possible to start from the divergence between different quantities (see the supplementary material [26]).
+We have considered the Kullback-Leibler divergence with respect to trajectories $\Gamma$, but it is also possible to start from the divergence between different quantities.
 Another direction of expansion is towards quantum systems.
@@
 This direction presents future challenges.
-Acknowledgements.
-This work was supported by JSPS KAKENHI Grant Number JP23K24915.
-Appendix A Derivation of the main result [Eq. (14)]
-Let us introduce the absolute random variable of $\Phi$, which is expressed by $|\Phi|$. The probability distribution of $|\Phi|$ is given by
+Appendix A Derivation of the main result [Eqs. (10) and (14)]
+Let us introduce the absolute value of the random variable $\Phi$, denoted by $|\Phi|$. The probability distribution of $|\Phi|$ is given by
 $P(|\Phi|=\phi)=\begin{cases}P(\Phi=\phi)+P(\Phi=-\phi)&\phi>0\\ P(\Phi=0)&\phi=0\end{cases}$ (22)
@@
 where we abbreviated $P(\phi)=P(\Phi=\phi)$.
-Using the monotonicity of Kullback-Leibler divergence [Eq. (5)] and the time-reversal property of $\Phi$ [Eq. (7)], we have
+Using the monotonicity of the Kullback-Leibler divergence [Eq. (5)] and the time-reversal property of $\Phi$ [Eq. (7)], we have
 $\Sigma$ $\geq D[P(\Phi)\|P(-\Phi)]$
@@
 Since the cross entropy term is no less than the entropy, it is non-negative, $C[P(\Phi),P(-\Phi)]\geq H[P(\Phi)]\geq 0$.
-An important observation is that $C[P(\Phi),P(-\Phi)]$ is even bounded from below by a positive term.
+The important observation is that $C[P(\Phi),P(-\Phi)]$ is even bounded from below by a positive term.
 We calculate $H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$ as follows:
@@
 $=\sum_{\phi}P(\phi)\ln\frac{P(-\phi)/P(\phi)}{1+P(-\phi)/P(\phi)}+P(0)\ln 2.$ (25)
-Let us consider the function $f(x)=\ln\frac{x}{1+x}$. Since $f(x)$ is concave for $x>0$, by using the Jensen inequality, the following relation holds:
+Let us consider the function
+$\mathfrak{f}(x)=\ln\frac{x}{1+x}.$ (26)
+Since $\mathfrak{f}(x)$ is concave for $x>0$, by using the Jensen inequality, the following relation holds:
 $\sum_{\phi}P(\phi)\ln\frac{P(-\phi)/P(\phi)}{1+P(-\phi)/P(\phi)}$
@@
 The Shannon entropy is defined by
-[…]
+$H[P(\Phi)]=-\int_{-\infty}^{\infty}d\phi\;P(\phi)\ln P(\phi).$ (29)
 For the continuous case, the probability density $P(|\Phi|)$ is defined by
-$P(|\Phi|=\phi)=P(\Phi=\phi)+P(\Phi=-\phi).$ (25)
-We do not consider $P(|\Phi|=0)$, because such measure is 00 for the smooth probability density.
+$P(|\Phi|=\phi)=P(\Phi=\phi)+P(\Phi=-\phi)\;\;(\phi\geq 0).$ (30)
+We do not consider $P(|\Phi|=0)$, because this set has measure zero for the smooth probability density.
 When $P(\Phi)$ includes the contribution of the delta function in $\Phi=0$, this is not the case.
@@
 $H[P(|\Phi|)]+\ln 2$ $\leq C[P(\Phi),P(-\Phi)].$ (31)
-Appendix B Symmetry entropy and Jensen-Shannon divergence
-The Jensen-Shannon divergence is defined by
-$\mathrm{JS}[P(X)\|P(Y)]$ $\equiv\frac{1}{2}D\left[P(X)\|\frac{P(X)+P(Y)}{2}\right]$
-$+\frac{1}{2}D\left[P(Y)\|\frac{P(X)+P(Y)}{2}\right],$ (27)
+Appendix B Proof of Eqs. (9) and (13)
+In this section, we prove Eqs. (9) and (13). Here, we show the relation for $P(\Phi=0)\geq 0$. The relation that we want to show is given by
+$0\leq\Lambda[P(\Phi)]\leq[1-P(\Phi=0)]\ln 2.$ (32)
+The first inequality part corresponds to $H[P(|\Phi|)]\leq H[P(\Phi)]$. Specifically, $H[P(\Phi)]-H[P(|\Phi|)]$ is
+$H[P(\Phi)]-H[P(|\Phi|)]$
+$=\sum_{{\phi>0}}[P(\phi)+P(-\phi)]\ln\left[P(\phi)+P(-\phi)\right]$
+${+P(0)\ln P(0)}-\sum_{\phi}P(\phi)\ln P(\phi)$
+$=\sum_{\phi>0}\Bigl[[P(\phi)+P(-\phi)]\ln\left[P(\phi)+P(-\phi)\right]$
+$-P(\phi)\ln P(\phi)-P(-\phi)\ln P(-\phi)\Bigr]\geq 0,$ (33)
+where we used $(a+b)\ln(a+b)-a\ln a-b\ln b>0$ for $a>0$ and $b>0$ in the last line.
+$(a+b)\ln(a+b)-a\ln a-b\ln b=0$ if and only if $a=0$ or $b=0$.
+This indicates that Eq.
+(33) saturates if and only if $P(\phi)=0$ or $P(-\phi)=0$ for all $\phi$ except $P(0)$.
+Next, we prove the second part of the inequality of Eq. (32), which can be done following the same approach as in Appendix A. Using Eq. (23), we have
+$H[P(|\Phi|)]-H[P(\Phi)]$
+$=-\sum_{\phi}P(\phi)\ln\left(1+\frac{P(-\phi)}{P(\phi)}\right)+P(0)\ln 2$
+$\geq-\ln 2+P(0)\ln 2,$ (34)
+where we again used the Jensen inequality. Equation (34) proves the second inequality part of Eq. (32). Equality in Eq. (34) holds if and only if $P(-\phi)/P(\phi)=1$.
+By adopting the formulation based on the Jensen-Shannon divergence [Eq.
+(11)], the second part of the inequality in Eq.
+(32) can be directly derived from Eq.
+(36).
+However, note that the first inequality part of Eq.
+(32) cannot be obtained from Eq.
+(36).
+Appendix C Symmetry entropy and Jensen-Shannon divergence
+We derive Eq. (11). Although we assume $P(\Phi=0)=0$ in Eq. (11), the derivation below also applies to the general case, including $P(\Phi=0)>0$. The Jensen-Shannon divergence is defined by
+$\mathrm{JS}[P(X)\|P(Y)]$ $\equiv\frac{1}{2}D\left[P(X)\middle\|\frac{P(X)+P(Y)}{2}\right]$
+$+\frac{1}{2}D\left[P(Y)\middle\|\frac{P(X)+P(Y)}{2}\right],$ (35)
 which satisfies
-$0\leq\mathrm{JS}[P(X)\|P(Y)]\leq\ln 2$ (28)
+$0\leq\mathrm{JS}[P(X)\|P(Y)]\leq\ln 2.$ (36)
 Using the Jensen-Shannon divergence, we have
@@
 where we used Eq. (23). Equation (37) is Eq. (11) in the main text.
-Appendix C Proof of Eq. (9)
-In this section, we prove Eq. (9). Here, we show the relation for $P(\Phi=0)\geq 0$. The relation which we want to show is given by
-$0\leq\Lambda[P(\Phi)]\leq[1-P(\Phi=0)]\ln 2$ (30)
-The first inequality part corresponds to $H[P(|\Phi|)]\leq H[P(\Phi)]$. For random variables $X$ and $Y$, the following relation holds:
-$H[P(X),P(Y)]$ $=H[P(X)|P(Y)]+H[P(Y)]$
-$=H[P(Y)|P(X)]+H[P(X)],$ (31)
-where $H[P(Y)|P(X)]$ is the conditional entropy:
-$H[P(Y)|P(X)]$
-$\equiv-\sum_{x,y}P(X=x,Y=y)\ln P(Y=y|X=x).$ (32)
-Substituting $X=\Phi$ and $Y=|\Phi|$, we obtain
-$H[P(|\Phi|)|P(\Phi)]+H[P(\Phi)]$ $=H[P(\Phi)]$
-$=H[P(\Phi)|P(|\Phi|)]+H[P(|\Phi|)]$
-$\geq H[P(|\Phi|)],$ (33)
-where we used $H[P(|\Phi|)|P(\Phi)]=0$.
-Equation (33) proves the first inequality part of Eq.
+Appendix D Equality condition
+We consider the equality condition of the thermodynamic entropic uncertainty relation [Eqs.
+(10) and (14)].
+The derivation of the thermodynamic entropic uncertainty relation employs two key inequalities:
+  1. 1.
+    Monotonicity of the Kullback-Leibler divergence under coarse graining.
+  2. 2.
+    The Jensen inequality applied to a strictly concave function.
+To saturate the thermodynamic entropic uncertainty relation, both inequalities must be saturated simultaneously. The distribution $P(\phi)$ is defined by
+$P(\phi)=\sum_{\Gamma}\delta\bigl(\Phi(\Gamma)-\phi\bigr)\mathcal{P}(\Gamma).$ (38)
+Then the monotonicity states
+$D\bigl[\mathcal{P}(\Gamma)\|\mathcal{P}(\Gamma^{\dagger})\bigr]\geq D\bigl[P(\phi)\|P(-\phi)\bigr].$ (39)
+Equality holds if and only if the log-ratio depends on $\Gamma$ solely through $\Phi(\Gamma)$, i.e., there exists a function $h$ such that
+$\frac{\mathcal{P}(\Gamma)}{\mathcal{P}(\Gamma^{\dagger})}\;=\;h\!\bigl(\Phi(\Gamma)\bigr).$ (40)
+There are two scenarios in which Eq. (40) is satisfied:
+Case 1 (Equilibrium)
+$\mathcal{P}(\Gamma)=\mathcal{P}(\Gamma^{\dagger})$ for all $\Gamma$.
+Case 2 (Bijectivity)
+$\Phi(\Gamma)$ is a bijective function.
+Next, we consider the Jensen inequality part.
+We applied the Jensen inequality to the strictly concave function $\mathfrak{f}(x)$ defined in Eq.
 (26).
-Note that, for the continuous case, this proof does not work, as the continuous conditional entropy $H[P(\Phi)|P(|\Phi|)]$ may take negative values.
-However, we can show that the first inequality part also holds for the continuous case by a direct calculation.
-Specifically, $H[P(\Phi)]-H[P(|\Phi|)]$ is
-$H[P(\Phi)]-H[P(|\Phi|)]$
-[…]
-[…]
-$=\int_{0}^{\infty}d\phi\,\Bigl{[}[P(\phi)+P(-\phi)]\ln\left[P(\phi)+P(-\phi)\right]$
-$-P(\phi)\ln P(\phi)-P(-\phi)\ln P(-\phi)\Bigr{]}.$ (34)
-Because $(a+b)\ln(a+b)-a\ln a-b\ln b>0$ for $a>0$ and $b>0$, $H[P(\Phi)]-H[P(|\Phi|)]\geq 0$ is proved.
-Next, we prove the second part of the inequality of Eq. (30), which can be done following the same approach as in Appendix A. Using Eq. (19), we have
-$H[P(|\Phi|)]-H[P(\Phi)]$
-$=-\sum_{\phi}P(\phi)\ln\left(1+\frac{P(-\phi)}{P(\phi)}\right)+P(0)\ln 2$
-$\geq-\ln 2+P(0)\ln 2,$ (35)
-where we again used the Jensen inequality. Equation (35) proves the second inequality part of Eq. (30).
-By adopting the formulation based on the Jensen-Shannon divergence [Eq.
-(11)], the second part of the inequality in Eq.
-(30) can be directly derived from Eq.
+The equality of the Jensen inequality is saturated if and only if the following condition is satisfied:
+$\frac{P(\phi)}{P(-\phi)}\;=\;\text{constant}.$ (41)
+Since Eq. (41) should hold for all $\phi$, Eq. (41) is equivalently stated as
+$P(\phi)=P(-\phi).$ (42)
+To saturate the thermodynamic entropic uncertainty relation, we require both (40) and (42). We now show that this forces equilibrium:
+  • In Case 1, $\mathcal{P}(\Gamma)=\mathcal{P}(\Gamma^{\dagger})$ already defines the equilibrium. Moreover, equilibrium implies $P(\phi)=P(-\phi)$, so Eq. (42) is automatically satisfied.
+  • In Case 2, $\Phi$ is bijective, so Eq.
+(40) allows any $\mathcal{P}(\Gamma)$ expressed via $\Phi(\Gamma)$.
+But Eq.
+(42) then demands $\mathcal{P}(\Gamma)/\mathcal{P}(\Gamma^{\dagger})$ be constant, again forcing $\mathcal{P}(\Gamma)=\mathcal{P}(\Gamma^{\dagger})$.
+In summary, the thermodynamic entropic uncertainty relation [Eqs. (10) and (14)] is saturated if and only if the system is in equilibrium.
+Appendix E Diffusion decision model
+E.1 Model
+In the main text, we consider the diffusion decision model, where the evidence accumulation process is modeled by Eq.
 (17).
-However, note that the first inequality part of Eq.
-(30) cannot be obtained from Eq.
-(28).
+Let $P(x;t)$ be the probability density of $x$ at time $t$.
+The corresponding Fokker-Planck equation is
+$\frac{\partial P(x;t)}{\partial t}=-\frac{\partial}{\partial x}[fP(x;t)]+A\frac{\partial^{2}P(x;t)}{\partial x^{2}}.$ (43)
+The probability current of Eq. (43) is
+$J(x;t)\equiv fP(x;t)-A\frac{\partial P(x;t)}{\partial x}.$ (44)
+A trajectory $\Gamma$ corresponds to a realization of the Langevin dynamics within the time interval $0,\tau]$.
+In the main text, we define the observable $\Phi(\Gamma)$ for a decision-making process.
+The probability density of $y(t)$ is given by the Gaussian distribution [49]:
+$P(y;t)=\frac{1}{\sqrt{4\pi At}}\exp\left[-\frac{\left(y-ft\right)^{2}}{4At}\right].$ (45)
+E.2 Discrete case
+In the discrete model, the observable is defined in Eq. (18). From Eq. (45), the observable probability distribution at time $\tau$ is
+$P(\Phi=+1)$ $=\int_{\theta}^{\infty}P(y;t)dy=\frac{1}{2}\mathrm{erf}\left(\frac{ft-\theta}{2\sqrt{At}}\right)+\frac{1}{2},$ (46)
+$P(\Phi=0)$ $=\int_{-\theta}^{\theta}P(y;t)dy=-\frac{1}{2}\mathrm{erf}\left(\frac{ft-\theta}{2\sqrt{At}}\right)$
+$+\frac{1}{2}\mathrm{erf}\left(\frac{ft+\theta}{2\sqrt{At}}\right),$ (47)
+$P(\Phi=-1)$ $=\int_{-\infty}^{-\theta}P(y;t)dy=\frac{1}{2}-\frac{1}{2}\mathrm{erf}\left(\frac{ft+\theta}{2\sqrt{At}}\right).$ (48)
+The probabilities in Eqs. (46)–(48) are used to calculate $\Lambda[P(\Phi)]$.
+E.3 Continuous case
+The symmetry entropy is given by $\Lambda[P(\Phi)]=H[P(\Phi)]-H[P(|\Phi|)]$ [Eq. (8)]. For the observable given by Eq. (21), each of the terms in $\Lambda[P(\Phi)]$ is
+$H[P(\Phi)]$ $=-\int_{-\infty}^{\infty}dy\,P(y;\tau)\ln P(y;\tau),$ (49)
+$H[P(|\Phi|)]$ $=-\int_{0}^{\infty}dy\,\Bigl[\left[P(y;\tau)+P(-y;\tau)\right]$
+$\times\ln\left[P(y;\tau)+P(-y;\tau)\right]\Bigr],$ (50)
+where $P(y;t)$ is Eq. (45).
+Acknowledgements.
+This work was supported by JSPS KAKENHI Grant Number JP23K24915.
+DATA AVAILABILITY
+The data that support the findings of this article are openly available [50].